### v2

# Climate Science and Control Engineering: Insights, Parallels, and Connections

Abstract
Climate science is the multidisciplinary field that studies the Earth’s climate and its evolution—past, present, and future. This simultaneously includes studying the natural and human-caused factors that meaningfully influence and often adversely impact it. At the very core of climate science are indispensable climate models that predict future climate scenarios, inform policy decisions, and dictate how a country’s economy should change in light of the changing climate. Climate models capture a wide range of interacting dynamic processes via extremely complex ordinary and partial differential equations—processes that span varying spatiotemporal scales. To model these large-scale complex processes, state-of-the-art climate science leverages supercomputers, advanced simulations, statistical methods, and machine learning techniques to predict future climate.
An area of applied math and engineering that is rarely studied or applied in climate science is control engineering—the analytical study of dynamic systems. Given that climate systems are inherently dynamic, it is intuitive to analyze them within the framework of dynamic system science. This perspective has been underexplored in the literature. Motivated by recent control engineering initiatives, in this manuscript, we provide a tutorial that: (i) introduces the control engineering community to climate dynamics and modeling, including spatiotemporal scales and challenges in climate modeling; (ii) offers a fresh perspective on climate models from a control systems viewpoint; and (iii) explores the relevance and applicability of various advanced graph and network control-based approaches in building a physics-informed framework for learning, control and estimation in climate systems. We also present simple and then more complex climate models, depicting fundamental ideas and processes that are instrumental in building climate change projections.
This tutorial also builds parallels and observes connections between various contemporary problems at the forefront of climate science and their control theoretic counterparts. We specifically observe that an abundance of climate science problems can be linguistically reworded and mathematically framed as control theoretic ones. From state estimation, reachability analysis, and feedback control to model order reduction and system identification, the parallels between the two disciplines are delineated by describing the climate science equivalent of these controls concepts. This perhaps merits further investigation of some of these problems through a control lens. Grounded in control engineering, this tutorial in particular offers a new and robust foundation to potentially enhance strategic decision-making and support effective mitigation efforts. Finally, and as a disclaimer, we make no claims regarding whether this framework will be exceptionally useful to solve some climate science problems—the tutorial merely presents a fresh outlook on the problem while building reasonable analogies.
Index Terms:
Climate modeling, control engineering, dynamic systems, climate-driven decision making, climate mitigation strategies.
Acronyms
CMIP Coupled Model Intercomparison Project
EBM Energy Balance Model
ESM Earth System Model
GCM General Circulation Model
GHGs Greenhouse Gases
IAM Integrated Assessment Model
IPCC Intergovernmental Panel on Climate Change
NOAA National Oceanic and Atmospheric Administration
NWP Numerical Weather Prediction
ODE Ordinary Differential Equation
PDE Partial Differential Equation
RCM Regional Climate Model
I Introduction and Tutorial Objectives
Climate change is one of the most pressing challenges of our era, influencing ecosystems, economies, and societies on a global scale. Accurate climate modeling is essential for predicting future scenarios and informing strategic policies aimed at mitigation and adaptation. These models simulate the complex interactions between the Earth’s atmosphere, oceans, land, ice, and biosphere—capturing feedback mechanisms, nonlinear behaviors, and the influence of human activities and their impacts.
In fact, climate scientists and policymakers focus on projecting future temperature changes and understanding the role of greenhouse gases (GHGs), particularly carbon dioxide (CO2), in driving these changes. The elevated concentrations of GHGs increase the greenhouse effect, leading to global warming and associated climate disruptions. Accurate projections of temperature and GHGs levels are vital for assessing impacts of climate change such as sea-level rise, extreme weather events, and ecosystem alterations, thereby guiding effective policy decisions.
Despite significant advances in computational power, statistical and stochastic methods, and machine learning integration, control engineering principles are rarely presented as an alternative to study climate-related problems. Climate systems are at their core dynamic in nature, governed by underlying physical laws and exhibiting complex behaviors that evolve over time. However, the application of control engineering principles (which are fundamentally suited to analyzing and managing such dynamic systems) remains relatively underexplored in the field of climate science. The objectives of this tutorial are three-fold: (i) introduce control engineering researchers to climate science and modeling; (ii) produce a fresh outlook on climate problems; and (iii) build parallels between various known and well-studied climate science problems and their control theoretic counterparts.
Figure 1: Climate as a dynamic system. In this tutorial, we explain how the Earth’s climate can be viewed as a complex dynamical system with inputs, outputs, internal states, and external disturbances. We describe how the key components of the climate system—atmosphere, hydrosphere, cryosphere, biosphere, and land surface—interact and evolve over time. We cover the roles of natural and anthropogenic inputs (e.g., solar radiation, green infrastructure, and natural variability) and disturbances (e.g., emissions and deforestation), and how they influence the system’s trajectory. We delineate the different types and levels of climate models and highlight how those models map the climate inputs and disturbances to evolving states and observable outputs such as global average temperature, in the form of scenario-based simulations and projections. Throughout the tutorial, we highlight and examine these dynamic processes through a control-theoretic lens and discuss how control tools can provide structure for understanding, simulating, and potentially influencing climate dynamics.
Why Control Theory for Climate Science and the Tutorial’s Audience. This tutorial is introduced to begin addressing this challenge from our perspective as control researchers. Over the past decades, control theory has taken a proactive role in tackling real-life problems in dynamic systems, including infrastructure systems (e.g., energy systems [1] and transportation networks [2]), aerospace engineering [3], epidemiology [4], and medicine [5]. Control engineering offers a structured framework to analyze dynamic systems, optimize their performance, and design robust solutions for managing uncertainties while ensuring performance guarantee. By leveraging this perspective, we can expand our understanding of climate models and uncover new pathways for their utility and implications. Furthermore, the integration of advanced control-based approaches, such as graph theory and network control, holds significant promise for building physics-informed frameworks that can bridge the gap between modeling, control, and decision-making.
The readership of this tutorial are either control systems research who are looking for applications for some of their theoretic techniques, control engineers interested in climate modeling and mitigation strategies, or climate scientists who are interested in applying new techniques to solve either vintage climate problems or more stressing and modern ones.
Some Efforts in Controls for Climate. While some efforts have been made to apply optimization and control techniques to climate-related challenges, they have largely focused on specific goals or methodologies distinct from the approach presented here. For example, a group of researchers has examined weather and climate systems as dynamic systems influenced by human activity, with a particular emphasis on aerosol injections. In the studies [6, 7], the authors formulated optimal control problems to explore deliberate weather and climate modification. The authors modeled the atmosphere as a closed-loop dynamical system, where state variables included atmospheric parameters such as temperature and pressure, while control variables represented human interventions like aerosol injections. Their objective was to modify natural atmospheric processes to achieve desired weather outcomes, framed as a dynamic optimization problem. In their most recent work [8], they applied optimal control theory to develop strategies for stratospheric aerosol injections, aiming to stabilize global climate temperatures by adjusting aerosol emission rates.
In addition, the study [9] has regarded the global climate system as a controlled dynamic system, with controls corresponding to economic activities causing emissions of GHGs. These efforts have applied nonlinear optimization techniques to solve dynamic control problems, aiming to maximize economic outcomes while adhering to environmental constraints. Similarly, the work in [10] provides a tutorial on the dynamic integrated model of climate and economy model, which employs discrete-time control to estimate optimal pathways for balancing GHGs emissions and economic development.
The study [11] takes a related approach by modeling the dynamic relationship between CO2 emissions and investments in reforestation and clean technology, highlighting the importance of efficiently allocating resources to achieve emission reduction targets such as those outlined in the Kyoto Protocol. Another earlier noteworthy effort in [12], has employed a two-stage optimization process to develop optimal CO2 emission reduction strategies, demonstrating sensitivity to regional and global environmental targets.
Recent Control Engineering Initiatives. The work in [13] explores the application of control theory to systematically guide climate modification efforts. The study demonstrates how control analysis can inform the design of climate regulators through simple model applications and acknowledges the potential of applying control theory to climate regulation pathways. The magazine paper [14] has highlighted the potential for the control systems community to contribute to tackling challenges in climate change for specific applications such as transportation infrastructure, electric power grids, and agriculture. For instance, they outline, at a conceptual level, how control principles could support the integration of renewable energy into power systems, improve transportation system efficiency, and enhance decision-making frameworks in climate-aware agriculture, but without presenting any mathematical models, formulations, or methodological details. That is, this study primarily serves as a call to action by identifying broad opportunity areas rather than developing specific actual methods or techniques for implementation. This tutorial takes motivation from [14].
These aforementioned studies either focus on specific optimization techniques for particular problems or abstractly discuss opportunities without detailing practical implications or actionable methods. In contrast, this manuscript aims to address this interdisciplinary gap by providing a comprehensive tutorial that explores the intersection of climate science and control engineering. This tutorial covers the climate dynamics and the practical control-theoretic techniques applicable to such dynamics. By doing so, we hope to foster a deeper understanding of the potential contributions that control theory can make to tackling climate-driven challenges and inspire innovative research in this critical area.
Figure 2: Projected global mean surface temperature changes relative to 1850–1900 levels under five emissions scenarios: very low emissions (SSP1-1.9), low emissions (SSP1-2.6), midlevel emissions (SSP2-4.5), high emissions (SSP3-7.0), and very high emissions (SSP5-8.5), illustrating the impact of different policy and socioeconomic pathways on future warming. Shading indicates 5% to 95% confidence intervals. Figure adapted from the IPCC Sixth Assessment Report (AR6) [15]. This tutorial shows how these (now culturally mainstream) figures are generated via dynamic models in climate science.
To offer a conceptual preview, Fig. 2 illustrates how global temperature trajectories are influenced by alternative policy and emission scenarios as published in a report aimed at policymakers [15]—more details on this report are included in the following sections. These projections, developed through complex dynamical models of the climate system, exemplify the types of problems addressed in this tutorial: how to model, predict, and control the evolution of climate variables. Throughout the following sections, we build the foundations necessary to understand how such projections are constructed, and how control engineering perspectives can contribute to designing effective climate interventions. We now proceed to delineate the main objectives and broader implications of this tutorial.
Tab. I: Climate Processes and Modeling List of References.
Topics Overview References
Atmospheric Processes Foundations of atmospheric physics and their mathematical representations [16, 17, 18, 19]
Ocean Processes Physical mechanisms and processes governing the ocean [20, 21, 22, 23, 24, 25]
Ocean Processes Modeling Introduction to the physical, mathematical, and numerical foundations of computer models used to understand and predict the global ocean climate system [26]
Atmosphere and Ocean Atmospheric and ocean processes and interactions [27, 28, 29, 30, 31, 32, 33, 34, 35]
Ocean and Climate Interactions between oceanic processes and the climate system, highlighting the oceans’ integral role in climate regulation [36]
Atmosphere and Climate Analysis and prediction of atmosphere processes and their feedbacks with the other systems spheres [37, 38]
Atmosphere/Temperature Variability Examines atmospheric temperature variability from a dynamical perspective and explores how it responds to climate change [39]
Cryosphere Processes Fundamentals and processes of ice sheets, glacier, and frozen water parts [40, 41, 42, 43, 44, 45]
Hydrosphere Focus on the physics and dynamics of all the water on earth — some references include the glaciers in this sphere [46, 40, 47]
Biosphere Overview of the evolution of the living organisms on Earth and their interactions with their environment and other spheres [48, 49, 50, 51, 52, 53]
Land Surface / Lithosphere Components and geodynamic processes of the lithosphere [54, 55, 56]
Climate Dynamics and Processes Climate components, process, interactions [57, 58, 59, 60, 61, 62]
Climate Models Climate models evolution, hierarchy, processes, comparisons, and limitations [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]
Regional Climate Models Focus on evolution and development of regional climate models [78, 79, 80, 81, 82, 83, 84, 85]
Climate Change Encompassing topics such as the physical science basis, observed impacts, mitigation strategies, policy frameworks, and adaptation measures [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104]
IPCC Assessment Reports IPCC has released six assessment reports on climate change, its causes, potential impacts, and response options [105, 106, 107, 108, 109, 15]
Tutorial Objectives and Broader Implications. This tutorial aims to build a foundational framework that connects the principles of control engineering with the complexities of climate science by:
  • Introducing Climate Dynamics and Models to the Control Community: We aim to provide control engineers and scientists with a comprehensive understanding of climate systems, highlighting their dynamic nature and the physical processes they encompass. We shed the light on how these dynamics can be conceptualized as controllable systems with multiple inputs, states, and outputs. We also survey the literature on the state-of-the-art in climate models, discussing their various scales included processes, limitations, and challenges. (Sections II and III)
  • Control-Oriented Representations of Climate Models: We introduce and present how climate models can be interpreted as control-oriented dynamical representations with the temporal evolution of variables/parameters as states, and system control inputs (e.g., in some cases, policies or human influence), and measured outputs of the systems corresponding to observable climate indicators. This structured, analytical approach enables the application of control and network-based techniques to climate systems. Such methodologies can enhance accurate modeling and prediction, determine optimal mitigation strategies, facilitate systematic analysis, and integrate and quantify uncertainties. (Sections IV, V, and VI)
  • Control-Theoretic Insights and Applications: We survey and demonstrate the relevance, potential, and applicability of control-theoretic and network-based techniques that can be applied to the control-oriented representations of climate dynamics. These techniques are further explored for their implications and outcomes in developing physics-informed frameworks for climate learning, control, and estimation. Leveraging such frameworks supports decision-making processes and drives mitigation efforts. Additionally, we identify potential research directions and practical applications that can benefit from this interdisciplinary approach. (Sections V, VI, and VII)
  • Drawing Parallels Between Climate Science Challenges and Control Engineering Concepts: We highlight conceptual and structural similarities between problems in climate science and control engineering. These include, for instance, the role of observability in inferring unmeasured climate states from sparse data. We also discuss the resemblance between climate model calibration and system identification, the parallels between uncertainty quantification in climate projections and robust control under model uncertainty, and the interpretation of mitigation policies as constrained optimal control problems. We note that although we make these observations in this manuscript, the aforementioned problems have not been studied through the lens of control engineering. By outlining these connections, we aim to foster a shared vocabulary and problem structure that can bridge the two fields and promote new interdisciplinary strategies. (Sections V, VI, and VII)
It is important to emphasize that the purpose of this tutorial is not to teach climate scientists control theory, nor to suggest that existing climate science lacks methodological depth. Rather, this tutorial is intended for the control engineering community, aiming to introduce climate systems as a rich, complex, and practically relevant domain where control-theoretic tools can be meaningfully applied. Climate models provide natural testbeds for developing and evaluating methods in large-scale dynamical systems, estimation under uncertainty, and robust decision-making—all areas where control theory has much to contribute. We simply describe what control systems theory can potentially deliver for climate research.
The structure of the tutorial is designed to support this goal and the aforementioned objectives: it begins with the basics and foundations of climate science and dynamics (Section II). Section III explains the evolution and classification of climate models. Section IV discusses the distinction between weather and climate and how these concepts can be interpreted within the control-oriented framework introduced in this tutorial. Section V highlights the dynamic nature of climate models from a control-theoretic perspective and poses questions that can be tackled using control methods. This is followed by Section VII, which examines climate dynamics through the lens of control engineering, offering interpretations and insights. In particular, this section draws explicit parallels between challenges in climate science and their counterparts in control theory. Section VIII concludes the tutorial with its summary and potential next steps and recommendations for the interested researchers.
The information presented in this manuscript on climate modeling and associated processes is based on the references listed in Tab. I. This compilation serves as a resource for researchers seeking foundational and advanced knowledge in the field. The table encompasses a range of topics, including atmospheric dynamics, ocean circulation, coupling climate dynamics, climate modeling, and climate change risks, providing a comprehensive overview of the literature that has shaped our current understanding of climate systems.
Tab. II: Summary of Climate System Components, Examples of their Governing Processes, and Modeling Characteristics.
Component Processes Examples & Modeled States Time Scale Spatial Scale Numerical Representation
Atmosphere Large-scale circulation, states: wind velocity components, pressure, temperature, specific humidity Days to weeks 100–1000 km Nonlinear PDEs: momentum, mass, and energy equations for a rotating, stratified atmosphere; vertical motion simplified using hydrostatic balance
Convection and turbulence, states: vertical velocity, heat and moisture fluxes Minutes to hours 1–10 km High-resolution nonlinear PDEs (Navier-Stokes + thermodynamics)—often parameterized
Radiative transfer, states: radiative fluxes (shortwave and longwave) Seconds to hours 1D vertical (column-wise), applied across global grid Radiative transfer integro-differential equations, often with semi-empirical approximations to account for absorption, scattering, and emission
Hydrosphere Ocean circulation (thermohaline, wind-driven), states: ocean currents, ocean temperature, salinity, sea surface height Months to millennia 10–1000 km Nonlinear PDEs: conservation of momentum, mass, temperature, and salinity under Earth’s rotation; solved using layered 3D ocean models
Surface wave dynamics, states: wave height, frequency spectrum, direction Seconds to hours 1-100 m Nonlinear PDEs (wave action balance equation) with algebraic source/sink terms (e.g., wind input, dissipation)
Cryosphere Ice sheet flow, states: ice thickness, velocity, basal temperature Years to millennia 10–1000 km Nonlinear PDEs for slow viscous flow, using shallow ice and shallow shelf approximations to simulate spreading, sliding, and thinning of ice
Snow cover and melt dynamics, states: snow depth, snow temperature, liquid water content Days to seasons 1D vertical (column wise) over a horizontal grid of scale 1–100 km 1D ODEs/PDEs describing snow temperature and melt based on energy balance; includes radiation, conduction, and phase change processes
Land surface Soil moisture and heat transport, states: soil moisture content, soil temperature profile Hours to seasons 1D vertical (column wise) over a horizontal grid of scale 0.1–10 km Coupled 1D PDEs for water and heat transport in the soil column (Richards’ equation and heat conduction)
Surface-atmosphere exchange (energy, momentum), states: surface temperature, latent and sensible heat fluxes, momentum flux Minutes to hours 1-10 m when modeled locally & grid-cell scale of the model when parameterized Algebraic flux formulations and empirical parameterizations for heat, moisture, and momentum exchange; sometimes coupled with simple ODEs (e.g., surface temperature evolution)
Biosphere Vegetation dynamics and photosynthesis, states: biomass, carbon assimilation rate Seasonal to decades 1–1000 km Coupled ODEs representing carbon uptake, biomass allocation, and plant response to environment
Biogeochemical cycling (carbon, nitrogen), states: carbon and nitrogen pools (e.g., soil organic carbon, plant carbon), nutrient fluxes Years to centuries Regional to global Coupled ODEs for elemental stocks and flows (e.g., carbon pools), with algebraic constraints for conservation and stoichiometry
Figure 3: Climate system’s main five components.
II Birds’ Eye View of Climate Models: From Simple Physics to the Discipline of Earth Science
This section provides a high-level overview of the climate system and its core components, setting the stage for how simple physical principles scale up to form the basis of Earth system science defined as the field that studies the Earth as one integrated system, where the interactions between its various components are examined and studied. These components are comprised of five main ones of the climate system (Fig. 3). A brief introduction to each is given next, while Tab. II summarizes the key physical and modeling characteristics of each. The table outlines the typical time scales (ranging from minutes to millennia), spatial scales (from meters to thousands of kilometers), and types of mathematical representations used in numerical climate models. As observed, each of these components is governed by fundamental conservation laws (e.g., of mass, momentum, and energy), with majority typically expressed through ordinary differential equations (ODEs) or partial differential equations (PDEs). These range from 1D vertical ODEs in snow and soil models, to high-dimensional nonlinear PDEs in atmosphere and ocean dynamics. This framing not only highlights the physics-based structure of climate models but also underscores their direct relevance to the principles of dynamic systems modeling and analysis.
The five main climate components: 1. Atmosphere: The layer of gases surrounding Earth, responsible for weather patterns, climate regulation, and the greenhouse effect. 2. Hydrosphere: All water on Earth, including oceans, lakes, rivers, and groundwater, play an important role in heat storage and transport. 3. Cryosphere: Frozen water on Earth, such as glaciers, ice sheets, and permafrost, affecting sea levels and Earth’s energy balance. Note that, some scientific disciplines and studies consider the cryosphere as part of the hydrosphere. 4. Land Surface: The solid Earth, including landforms, soils, and surface processes, which influence climate through heat exchange, water storage, and interactions with vegetation. 5. Biosphere: All living organisms, from microbes to forests, that interact with the climate system through carbon, water, and energy cycles.
Figure 4: Key transitional milestones in the history of climate science.
Modeling these components and their interactions is an intensively complex task. Yet, the small advances across many scientific areas have shaped the Earth science discipline and climate modeling. This progression dates back to at least the seventeenth century, beginning with the study of atmospheric science. The early efforts treat phenomena as stand-alone dynamics, translating them into mathematical formulations based on physical principles or observational data. Over time, researchers has begun studying interactions and coupling between these phenomena, which has led to the development of the early climate models. With advances in numerical simulations and the information revolution, modern climate models have become increasingly complex and arguably more accurate.
A high-level overview of the key transitional milestones in the history of climate science is illustrated in Fig. 4. These milestones have shaped climate and Earth science into one of the most widely discussed topics in theoretical research, academia, and policy. We present a brief history of climate science in what follows.
  1. 1.
    Early Theoretical Foundations:
    • 1824: The French physicist Joseph Fourier has proposed what we know now as greenhouse effect on the atmosphere’s heat and accordingly temperature.
    • 1896: The Swedish chemist Svante Arrhenius has manually quantified the impact of atmospheric CO2 on Earth’s temperature, laying the groundwork for future climate studies [110].
  2. 2.
    Early Climate Models and Advent of Numerical Weather Prediction:
    • 1922: The British scientist Lewis Fry Richardson has pioneered numerical weather prediction by attempting to calculate weather changes using mathematical equations, though his methods were limited by the computational tools of his time [111].
    • 1950: The first successful computerized weather forecast is generated using the Electronic Numerical Integrator and Computer (ENIAC) computer, marking a significant leap in meteorological science [112].
    • 1956: The Meteorologist Norman Phillips has developed a two-layer atmospheric model capable of simulating general atmospheric circulation, representing a foundational climate model [113, 114].
    • 1967: Syukuro Manabe and Richard Wetherald have introduced a one-dimensional atmospheric radiative-convective model, demonstrating that doubling CO2 levels could lead to a 2C increase in global temperatures. These results have been marked as a landmark finding in climate science, providing early evidence of human-induced climate change [115].
    • 1969: Syukuro Manabe and Kirk Bryan have introduced the first attempt of coupling the atmosphere and ocean dynamics in one compact model [116].
    • 1980s: This decade marked the start of significant advancements in climate modeling, driven by increased computational power. These improvements have enabled more complex models that captured interactions between the different climate components such as atmosphere, oceans, and land surfaces, leading to the more accurate climate projections used today.
  3. 3.
    Public Climate Forecasts and Policy Impact:
    • 1988: NASA scientist James Hansen has testified before the US Congress, presenting model-based evidence of global warming, bringing climate modeling to the forefront of public policy discussions [117].
    • 1990: The Intergovernmental Panel on Climate Change (IPCC) has released its First Assessment Report, heavily relying on climate models to inform global policy on climate change [105]. Since then, IPCC has published six more assessment reports in total, with the most recent being the Sixth Assessment Report (AR6) from 2021–2023. The IPCC is now in its seventh assessment cycle, which began in July 2023, with the Seventh Assessment Report (AR7) expected by late 2029.
  4. 4.
    Integration of Machine Learning (ML) and Artificial Inelegance (AI):
    • 2010s: Researchers have begun exploring AI and ML to improve climate models, focusing on enhancing the simulations and predictions accuracy [118].
    • 2020s: AI-driven models have been emerging, offering faster and potentially more accurate weather and climate forecasts by analyzing extensive datasets and identifying complex patterns [119].
Figure 5: Replica of the chronological evolution of climate modeling from IPCC report summarized in [120], extended by explicitly including the anthroposphere which represents human activities and their influence on the climate system. Compared to [120], this figure also maps processes to their respective climate system components and uses thicker arrows to highlight the increasing contribution of human activity to climate change, now explicitly incorporated into climate modeling frameworks.
The Anthroposphere: Modeling Our Mark on the Planet. Another important element influencing the climate, beyond natural elements, is the anthroposphere. This term refers to parts of the Earth’s system that humans have created or modified for activities and habitats [121]. Such activities include urbanization, deforestation, and GHGs emissions, leading to land-use changes and alterations in the atmospheric radiation balance. In many climate models, these human-induced processes are incorporated as external forcings, treated as prescribed inputs to the system based on historical data or scenario-based projections. Fig. 6 and the extended schematic of climate modeling evolution emphasizes the growing influence of human activities on climate dynamics. As models become more complex, the anthroposphere is no longer treated as merely external, but as an active component interacting with other parts of the climate system, revealing the importance of incorporating societal drivers into Earth system modeling.
Tab. III: Overview of Prominent Climate Models: Types, Dynamics, Applications, and Limitations.
Model Type Resolution & Dynamics Included Main Applications Limitations
Energy Balance Models (EBMs) Simplified representations focusing on the balance between incoming solar radiation and outgoing terrestrial radiation Estimating global mean temperature responses to changes in GHGs concentrations Lack spatial resolution; fails to capture regional climate variations or complex feedback mechanisms
General Circulation Models (GCMs) Comprehensive models simulating atmospheric and oceanic circulation, along with their interactions with land and ice Projecting global climate change scenarios; studying large-scale climate phenomena Computationally intensive; diner-scale processes simplifications lead to predictions uncertainty
Regional Climate Models (RCMs) High-resolution models focusing on specific regions, incorporating local topography and land-use data, which are often not adequately represented in larger GCMs Assessing regional climate impacts; informing local adaptation strategies; climate-aware regional urban planning Dependent on boundary conditions from GCMs; access to detailed regional land use and topography data can be limited
Earth System Models (ESMs) Extend GCMs by including biogeochemical cycles, such as carbon and nitrogen cycles Studying interactions between climate and ecosystems; assessing feedback in the Earth system; policy implications by assessing climate change impacts on ecosystems and human societies Increased complexity adds to computational demands; uncertainties in biogeochemical processes and interactions
Integrated Assessment Models (IAMs) Combine climate models with socio-economic factors to assess policy impacts on climate change Evaluating mitigation and adaptation strategies; informing policy decisions Simplifications may overlook critical interactions; results can be sensitive to assumptions and value judgments of climate change impacts
III Climate Models: Evolution, Scales, and Applications
While early climate models evolve progressively, as explained in the previous section, the development of modern climate models has become a more complex process. However, it has also become more achievable with the computational tools available today. This section, we present what climate models are known for: their evolution, their relevance to the various climate components, and what kind of climate models are used for specific applications.
III-A Evolution and Classification of Climate Models
In Fig. 5, we present the chronological evolution of climate models as documented in the IPCC report [108] and summarized in [120], with an important addition: we highlight which processes belong to which of the five main climate system components. As illustrated, efforts have focused on individual processes till they are well-established before integrating them into unified models that incorporate multiple interacting components.
Climate models are classified based on several aspects, including but not limited to: (i) dimensionality, (ii) spatial and temporal resolution, and (iii) the processes and dynamics included. These aspects are interdependent, meaning that the choice of one affects the others. For example, in a two-dimensional space model, the included processes may be limited to large-scale atmospheric or oceanic circulations, excluding finer-scale turbulence or convection that happens at a coarser spatial scale. The selection of processes, in turn, influences the required temporal resolution, as certain dynamics demand finer time steps to ensure stability of numerical solution to PDEs and ODEs. Additionally, spatial resolution affects both the level of detail captured and the computational feasibility.
A generalized classification of climate models is introduced in Tab. III. This classification includes the most widely utilized and prominent models, including EBMs, GCMs, RCMs, ESMs, and IAMs (acronyms are defined in the table), and lists their applications and limitations. In addition to the models outlined in Tab. III, microclimate variations and localized climate impacts over specific areas can be assessed using Local Climate Models (LCMs). These models offer high-resolution simulations tailored to these areas and are developed by downscaling the RCMs surrounding them, thereby informing urban planning and localized risk assessments. On the other hand, an emerging expanded version of the ESMs is the Human Earth System Models (HESMs) that aim to include more detailed representations of human decision-making and social dynamics within the Earth system model.
In connection with the climate system’s main five components in Section II, we emphasize herein the different combination of these components in these models classified in Tab. III depending on their purpose and complexity. The EBMs typically represent the atmosphere and land surface in a highly simplified form, often through globally averaged equations, to capture large-scale temperature responses. GCMs explicitly simulate the atmosphere and ocean (hydrosphere), and often include simplified representations of the cryosphere, land surface, and biosphere. RCMs refine the spatial resolution of GCM outputs over specific regions, enabling more detailed representations of land surface heterogeneity and localized atmospheric processes. ESMs expand upon GCMs by incorporating interactive biosphere and biogeochemical cycles, such as carbon and nitrogen flows, which play an important role in representing long-term climate feedbacks. IAMs take a broader systems perspective by loosely coupling simplified representations of all five components with economic and policy modules to support decision-making. LCMs, derived from RCMs, focus on fine-scale atmospheric and land surface interactions relevant to urban microclimates. Finally, HESMs aim to capture the coupled dynamics between human systems and the climate, requiring deeper integration of land, biosphere, and atmospheric processes with socio-economic behavior. As shown in Tab. II, the modeling of each component involves different characteristic time and spatial scales, numerical approaches, and governing equations, which influence model structure and selection.
To that end, while this classification provides a structured overview of climate models levels and types, it is important to note that under each category, multiple models exist, each designed with different assumptions, computational frameworks, and objectives. These models differ in structural design, spatial resolution, parameterization schemes, and numerical methods used to solve the complex equations governing climate processes. Parameterization refers to the process of approximating small-scale processes that cannot be directly resolved, such as cloud formation and convection. The choice of a particular model depends on the specific research question, the scale of analysis, and the computational resources available.
For instance, in their Sixth Assessment Report (AR6), IPCC utilized outputs from the Coupled Model Intercomparison Project Phase 6 (CMIP6). This project encompasses simulations from approximately 100 distinct climate models developed by 49 different modeling groups worldwide. These models provide a comprehensive foundation for assessing future climate scenarios and understanding potential impacts. To enhance accessibility and facilitate flexible spatial and temporal analysis of climate data, the IPCC’s AR6 also introduced the Interactive Atlas [122]. This interactive tool allows users to explore regional and global climate projections derived from CMIP6 models, making complex climate information more accessible. Readers can leverage this platform to analyze climate projections, assess regional impacts, and gain deeper insights into climate variability and change.
Figure 6: Figures generated by the IPCC AR6 Atlas [122], showing projected changes in mean annual temperature (°C) relative to the 1850–1900 baseline under different emission scenarios leading to average global warming of $1.5^{\circ}$, $2^{\circ}$, $3^{\circ}$, and $4^{\circ}$. Solid areas indicate high agreement among CMIP6 models (34 models for the first three scenarios, 20 for the last), while hatched areas indicate low agreement.
An Example on Predicting and Computing Climate Projections. Here, we illustrate how climate projections—that have become part of the climate discourse—are computed via dynamic climate models. For example, Fig. 6 presents visualizations generated by the IPCC’s Interactive Atlas, displaying projected changes in mean annual temperature relative to the 1850–1900 baseline [^1] The 1850–1900 baseline is used by the IPCC as a proxy for pre-industrial conditions, as it represents the average temperature before significant industrialization and the associated increase in GHGs emissions. It is computed by averaging observed or reanalyzed surface temperatures over that period. While direct instrumental coverage was sparse, especially in the Southern Hemisphere, recent studies [123, 124] show that the baseline is accurate within about ±$\pm 0.1^{\circ}$, making it suitable for long-term comparisons and global warming targets. under four global warming levels: $1.5^{\circ}$, $2^{\circ}$, $3^{\circ}$, and $4^{\circ}$. These specific levels represent policy-relevant thresholds and scientific milestones in the context of climate change mitigation and adaptation. The $1.5^{\circ}$ and $2^{\circ}$ targets are directly tied to the Paris Agreement, [125]—an international treaty adopted under the United Nations Framework Convention on Climate Change (UNFCCC) that aims to limit global warming to well below 2C, with efforts to pursue a 1.5C target above pre-industrial levels. Each additional degree of warming brings increasingly severe consequences. That being said, the $3^{\circ}$ and $4^{\circ}$ levels serve as illustrative benchmarks for high-emission scenarios. These levels are also widely used in impact assessments across various domains, such as ocean science, water availability, energy systems, agriculture, and public health, where researchers estimate physical and societal impacts under different degrees of warming. This makes them a common language for linking model projections to real-world consequences.
These projections are derived from a multi-model ensemble of CMIP6 simulations, using 34 dynamic models for the first three scenarios and 20 for the fourth. The reduced number of models at the $4^{\circ}$ warming level is primarily due to the fact that not all modeling centers have submitted simulations for high-emission or extended scenarios, and some models were excluded due to missing output fields or inconsistencies that affect comparability. From a systems and dynamics perspective, each of these models solves a set of coupled nonlinear ODEs and PDEs that govern the evolution of climate variables such as temperature, wind, ocean currents, and moisture. These PDEs are discretized over spatial grids and time steps, often with horizontal resolutions ranging from tens to hundreds of kilometers, which means that generating a single simulation involves high computational cost and long runtimes. For example, simulating a century-long climate trajectory at high spatial and temporal resolution requires significant computational infrastructure, limiting how many runs some centers can produce, especially for extreme scenarios. The plots shown are constructed by spatially averaging the simulated surface temperature fields over time for each model grid, followed by computing the ensemble mean across all models included for a given warming level. Agreement and confidence are visually encoded: solid-shaded regions indicate strong agreement among the models (i.e., consistent sign of change in at least 80% of models), while hatched regions indicate areas of low confidence where models disagree on the direction of change or the signal-to-noise ratio is weak.
The Interactive Atlas offers additional features for other climate variables such as surface wind and precipitation, and enables users to generate regional time series, assess seasonal variability, and explore different emissions scenarios through an interactive interface. Tools such as AR6 Atlas and the CMIP6 simulations are a rich source that can be used as testbed for generating data that can be used to solve climate-relevant control engineering problems as emphasized in Section VII.
III-B How are Climate Models Useful?
Beyond predicting the adverse effects of climate change, climate models can be useful for a plethora of applications that also involve dynamic system models. Indeed, climate modeling is exploited as a preliminary tool across various sectors to provide the needed insights for real-time decision-making in infrastructure planning, resource management, disaster preparedness, and more. These models help predict long-term trends and short-term variations, allowing industries and governments to make informed choices that are a function of physics-based climate models.
One of the most significant drivers of climate modeling is that it serves as the foundation for understanding, predicting, and responding to climate change. These models help quantify the impact of GHGs emissions, assess long-term climate trends, and evaluate the potential consequences of different mitigation and adaptation strategies. The insights derived from climate models influence global policy frameworks, such as the Paris Agreement mentioned in the previous section, and guide regional and national climate action plans.
That is, climate models have broad applications across various domains, where projects, trends, variations, shifts, and changes in climate variables and processes inform these domains’ practices. The scope of these domains is wide, with varying levels of focus depending on the application and expected outcomes. Below, we outline some of these domains to give readers an overview and highlight the broader significance of these models beyond understanding the Earth and its climate. We also acknowledge the entities involved in investigating and addressing these areas, as well as those providing data and resources that support potential opportunities.
  • Infrastructure and Urban Planning: Cities rely on climate projections to build resilient infrastructure, develop flood control strategies, and plan sustainable transportation systems in response to shifting climate patterns. NOAA’s National Centers for Environmental Information (NCEI) [126] provide climate data that assist in planning resilient infrastructure and urban development to withstand changing climate conditions.
  • Power/Energy Systems and Integration of Renewables: Climate models help optimize renewable energy deployment by forecasting long-term wind, solar, and hydrological trends, ensuring efficient energy grid management under varying climate conditions. These models provide the needed data for integrating wind and solar farms, optimizing hydropower operations, and planning energy storage solutions to balance supply and demand. Several organizations leverage climate modeling for energy planning. For instance, the International Energy Agency (IEA) [127] uses climate projections to assess renewable energy potential and develop policies for sustainable energy transitions. The National Renewable Energy Laboratory (NREL) [128] integrates climate models into its Renewable Energy Integration studies, evaluating how weather and climate variability impact grid stability and energy reliability. Additionally, the European Centre for Medium-Range Weather Forecasts (ECMWF) provides high-resolution climate data that support the optimization of wind and solar energy production [129]. The APEC Climate Center (APCC) [130] conducts research on climate predictions, which supports the integration of renewable energy sources by forecasting climate variability. In short, dynamic climate models inform and occasionally impact other dynamic models in energy and power systems.
  • Agriculture and Food Security: Farmers and policymakers use climate models to anticipate drought risks, changes in growing seasons, and extreme weather events, enabling adaptive agricultural practices and food security planning. Climate Information Services (CIS) [131] offer forecasts and agrometeorological services, including tools that enable farmers to make informed decisions, enhancing food security.
  • Water Resource Management: Climate models inform reservoir operations, groundwater sustainability, and drought mitigation strategies, ensuring water availability for urban, industrial, and agricultural use. Through its Water Resources program, NOAA’s (National Oceanic and Atmospheric Administration) Climate Program Office (CPO) advances approaches to prepare for and respond to climate-related impacts on water resources, supporting research and the development of strategies to address climate variability [132]. Additionally, the International Water Management Institute (IWMI) integrates modeling, monitoring, and scenario planning for surface water and groundwater to help governments and partners plan and operate for climate change adaptation, further developing and deploying flood and drought monitoring and forecasting technologies [133]. Similarly, RTI Center for Water Resources utilizes advanced, information-based technology and solutions to translate data into actionable information, strengthening water security and resilience against increasing demand, water scarcity, extreme weather events, and hydrologic uncertainty [134]. In the United States, the Water Utility Climate Alliance (WUCA) emphasizes practical approaches to climate change adaptation, planning, and decision-making to ensure that water utilities and the communities they serve can thrive amidst emerging challenges. By leveraging collective utility experiences, WUCA develops leading practices in climate change adaptation and mitigation that are actionable, equitable, and serve as models for others [135].
  • Disaster Preparedness and Risk Management: Governments and humanitarian organizations rely on climate models to develop early warning systems for hazards such as hurricanes, heatwaves, and wildfires, improving disaster response strategies. Initiatives like those led by the United Nations Office for Disaster Risk Reduction (UNDRR) [136] are taken to promote comprehensive disaster and climate risk management as central to development planning, emphasizing the integration of risk-centered approaches into national adaptation plans.
Other dynamic system application domains that climate models inform and impact are traffic networks, air quality, wildfires prediction, and spread of diseases.
IV The Weather We Experience vs. The Climate We Predict
Having introduced the climate components and classified various climate models with their applications, we highlight in this section the fundamental difference between weather and climate before presenting some mathematical climate models in Section V.
The concepts of weather and climate are sometimes confused and used interchangeably. Weather refers to the short-term atmospheric conditions in a specific place at a specific time. It encompasses daily variations in temperature, humidity, precipitation, wind, and visibility, along with severe conditions including droughts, tornadoes, and hurricanes. These conditions can change within minutes or hours. In contrast, climate represents the long-term average of weather patterns over an extended period—months, years, or even centuries. As NOAA puts it: Climate is what you expect, weather is what you get.
From a control systems perspective, this distinction is analogous to analyzing a system’s short-term response (transients) versus its long-term behavior (average or steady-state response). While weather prediction involves tracking the immediate state trajectory of a highly nonlinear, chaotic system, climate modeling focuses on its statistical or average behavior over time making concepts related to infinite-horizon or asymptotic stability, reachability, and observability relevant for climate dynamics. This shift in focus from state evolution to long-term system characteristics is key for developing control-oriented strategies for climate adaptation and mitigation.
IV-A Entities Providing Weather Forecasts and Climate Projections
Several organizations worldwide are responsible for monitoring atmospheric conditions and providing weather forecasts, which influence daily activities, business operations, and decision-making. These organizations use various forecasting models—typically referred to as numerical weather prediction (NWP), differing in prediction timeframes and spatial resolutions. These agencies, among others, are part of the World Meteorological Organization (WMO), a specialized agency of the United Nations (UN). The WMO acts as a coordinating body for data sharing and best practices in global weather forecasting.
  • The National Weather Service (NWS) in the US: Federal agency working under the umbrella of NOAA. The NWS operates two main forecasting models: the North American Mesoscale (NAM), which regional numerical weather prediction for the North American Continent, and the Global Forecast System (GFS), a global numerical weather prediction system. For more details on these models, refer to their official website [137].
  • European Centre for Medium-Range Weather Forecasts (ECMWF): An intergovernmental organization supported by European nations, providing global weather predictions and data. ECMWF operates the Integrated Forecasting System (IFS) [129].
  • Met Office in the UK: The UK’s national weather service, delivering weather and climate-related services. The Met Office utilizes the Unified Model, a numerical weather prediction and climate modeling software suite [138].
  • Japan Meteorological Agency (JMA): Japan’s national weather service, which operates ensemble prediction systems for medium-range forecasts [139].
  • Bureau of Meteorology (BoM) in Australia: Uses the Australian Community Climate and Earth-System Simulator (ACCESS) for weather and climate modeling [140].
In addition to the specialized models employed by various meteorological agencies, the Weather Research and Forecasting (WRF) Model stands out as a versatile, open-source numerical weather prediction system. Developed collaboratively by U.S. agencies and research institutions, WRF serves both atmospheric research and operational forecasting needs. WRF is distinguished for its adaptability across different scales, advanced physics and numerical methods, and data assimilation capabilities—for more detailed information visit [141].
Why is this information about these meteorological agencies and their forecasting models important for the controls community? Many of these agencies provide open access to full or partial datasets of historical and real-time weather data for academic research and scientific applications. Additionally, analyzing the differences in their model structures, time-frames, and methodologies and linking them to control engineering dynamics and applications discussed in this tutorial creates opportunities for further research and practical implementations.
IV-B Relevance to Controls Community
In this section, we pose an important question that pertains to the potentially important role the control engineering community can play in answering some climate-related questions.
Are the control-theoretic insights and interpretations proposed in this tutorial applicable to weather forecast models? The general answer is yes, as these models are dynamical systems with close applications to the climate models. However, acknowledging the following key differences between climate and weather forecasting models to ensure that the application of control-theoretic insights is appropriately framed. Temporal Scales: Weather models focus on short-term forecasts (hours to weeks) whereas climate models simulate long-term trends over years to centuries. Required Conditions: Weather models rely on initial conditions, which define the system’s state at the start of a forecast. This accordingly highly impacts short-term accuracy. Climate models, in contrast, depend more on boundary conditions and external forcings, which set constraints like GHGs levels and ocean heat content to capture long-term trends. Resolution and Computational Complexity: Weather models do not include processes that do not significantly impact short-term forecasts, such as atmospheric chemistry and the carbon cycle. On the other hand, climate models consider these processes to capture long-term interactions between the atmosphere, oceans, land, ice and biosphere. Additionally, they focus on long-term averages and trends, often smoothing out high-frequency fluctuations that influence weather forecast.
Thus, while control-theoretic approaches are applicable, they must be tailored to the specific constraints and objectives of each model type. The next sections delve into actual dynamic models, highlighting the relationship between these models and the aforementioned forecasts and projections.
V Simple Dynamic System Modeling of Climate Processes
Climate processes evolve over time, highlighting the dynamical nature of the system, which is driven by complex interactions between climate components. Climate models are built on fundamental physics-based principles. This includes conservation of energy, mass, and momentum. After obtaining these models, the interactions between these components over space and time are studied, thereby capturing the variability and feedback mechanisms that shape Earth’s climate. In this section, we showcase this dynamical nature through one of the simplest energy balance climate models (EBM). EBMs serve as a starting point and a pedagogical tool to motivate the more complex representations of climate dynamics discussed in Section VI.
EBMs connect back to the broader modeling characteristics summarized in Tab. II and the classification outlined in Tab. III, by exemplifying how even a minimal model can represent key climate behaviors in a control-oriented form. Historically, such simplified EBMs played a foundational role in early climate studies, offering insight into climate sensitivity, radiative forcing, and the concept of planetary energy balance. These early models have laid the groundwork for the development of modern, high-resolution climate models.
V-A Dynamic Climate Models 101: The Unforced EBM
To introduce the control-oriented interpretation of climate models, we begin with a simple yet insightful example. The following EBM is a zero-dimensional model that simulates the change in the global mean surface temperature [^2] It may sound unintuitive, and frankly confusing, that the global mean temperature is not really an average or mean of anything. While the reader might think that this ODE models the average across different spatial $T_{i}(t)$ (each with different parameters $C_{i},S_{i},\alpha_{i},\ldots$, in reality the model just considers the average parameters into a single ODE with one state. In short, this simplification works by reducing all spatial variability into a single time-evolving state variable $T(t)$. This approach captures large-scale energy balance trends without needing spatial resolution. over time. This model is expressed as a first-order ODE as in Eq. (1):
$\leavevmode\nobreak\ C\frac{dT(t)}{dt}=S(1-\alpha)-\epsilon\sigma T(t)^{4}.$ (1)
[…]
This model describes how temperature $T(t)$ evolves continuously over time rather than giving a static or equilibrium value. The system responds to external conditions such as changes in solar radiation, albedo, or GHGs. The presence of $\frac{dT(t)}{dt}$ makes the model time-dependent, and the feedback term $\epsilon\sigma T^{4}$ introduces nonlinearity, which governs how the system stabilizes or amplifies—a more detailed discussion on this aspect is included later in this section.
From a control perspective, this model behaves as a nonlinear first-order system with a single state and no explicit control input, meaning that the system evolves on its own based on the physical parameters including solar radiation, albedo, and GHGs levels. The solution to this ODE gives the trajectory of global temperature over time, which can be analyzed numerically.
V-B Modeling Control Inputs and GHGs Emissions: Formulating a Forced EBM
To influence the evolution of $T(t)$ in (1), control inputs can be appended to the dynamics. In the context of this simple EBM (1), these control inputs include geoengineering techniques and policy levers, as they can modify the model structure and/or parameters. As for GHGs emissions, they are typically not included or modeled as control inputs in this framework. Instead of modeling GHGs directly via some other ODEs or even PDEs, we can alternatively model actions that influence emissions such as:
Carbon taxes or regulations $\rightarrow$ reduce emission rates. Renewable energy adoption $\rightarrow$ displaces fossil fuels. Carbon capture technology $\rightarrow$ removes CO2 from the atmosphere.
These policy or technological levers can be modeled as control inputs that affect the rate of emissions, which affects atmospheric CO2 concentration, which affects emissivity $\epsilon$, which affects the feedback term $\epsilon\sigma T^{4}$, which changes the temperature evolution.
Similarly, one of the geoengineering techniques that can be treated as control input is the stratospheric aerosol injection. Stratospheric aerosol injection is a geoengineering technique that involves releasing tiny reflective particles (usually sulfate aerosols) into the stratosphere (about 10–50 km above the Earth’s surface) to reflect a portion of incoming solar radiation back into space. An interesting control engineering problem that can be posed here is how would introducing these aerosols injections impact the mean global temperature dynamics? Conceptually, the below simple flow diagram illustrates the idea.
Stratospheric aerosol injection $\rightarrow$ reflects sunlight $\rightarrow$ increases albedo $\alpha$ in (1) $\rightarrow$ less energy absorbed $\rightarrow$ temperature $T(t)$ drops in (1)
In particular, we modify the EBM in (1) into (2) to include ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}$ as the control input, representing the effective albedo increase from aerosol injection. ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}$, bounded based on technical and environmental limits. By introducing this control input, the term $S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))$ decreases, which means less energy absorbed, and thus cooling.
$\leavevmode\nobreak\ C\frac{dT(t)}{dt}=S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))-\epsilon\sigma T(t)^{4}.$ (2)
So the broad idea here is as follows. By default, the model (1) is an open-loop one but it can be extended to a closed-loop system (2) by introducing such inputs. In addition to incorporating deliberate control inputs such as aerosol injection through ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}$, external disturbances can also be introduced to capture the effect of GHG emissions. A common approach is to treat changes in emissivity $\epsilon$ as an uncertain and possibly time-varying quantity. Rising GHG concentrations reduce Earth’s ability to emit energy back into space, which is effectively modeled by a decrease in $\epsilon$.
Instead of solving an additional ODE for atmospheric CO2 concentration, we capture its impact on temperature dynamics by modifying $\epsilon$ as a time-varying function $\epsilon(t)$, or by modeling its deviation from a nominal value using an additive disturbance term ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)}$. This leads to a disturbed version of the energy balance model:
$\leavevmode\nobreak\ C\frac{dT(t)}{dt}=S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))-(\epsilon+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)})\sigma T(t)^{4},$ (3)
where ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}$ is the control input, representing interventions such as stratospheric aerosol injection. However, variable ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)}$ is an external disturbance capturing the impact of increased GHG concentrations on emissivity.
If no explicit control is applied (i.e., $u(t)=0$) and emissions rise unchecked (i.e., $w(t)<0$), the system experiences increased warming due to enhanced radiative trapping. Conversely, effective control ($u(t)>0$) and stable or reduced emissions ($w(t)\approx 0$) can lead to stabilization or even gradual cooling.
Figure 7: Illustrative simulation of global mean temperature evolution under three scenarios using a zero-dimensional EBM: an unforced baseline, a scenario with control-based cooling intervention, and a scenario with external radiative forcing (disturbance), e.g., GHGs. Each trajectory is simulated under process-level uncertainty modeled as multiplicative noise applied to the system dynamics. Shaded regions represent the 5th to 95th envelopes over 100 Monte Carlo simulations.
To introduce the idea of framing the climate dynamics in (1) in a control-theoretic framework, Fig. 7 presents an illustrative simulation that showcases how the global mean surface temperature evolves over time in three distinct scenarios: the unforced model representing the baseline climate system, a disturbed model where an external forcing (e.g., rising GHG emissions) reduces Earth’s emissivity, and a controlled model where an artificial albedo increase is used to counteract warming. This example highlights a core idea from control theory: external disturbances such as increased emissions can push the system away from its nominal behavior, while appropriately designed control inputs can counterbalance these effects and reach desirable states. In this specific case, the control input mimics geoengineering strategies (e.g., stratospheric aerosol injection) that increase planetary albedo, leading to lower absorbed solar radiation and thus reduced surface temperatures.
To reflect uncertainty in physical processes and imperfect model knowledge, we augment the system with two sources of uncertainty: (i) process-level stochastic noise affecting the energy balance dynamics, and (ii) random perturbations in key parameters, specifically the albedo $\alpha$ and the emissivity $\epsilon$. A total of 100 Monte Carlo simulations are performed for each scenario (unforced, controlled, and disturbed), with each realization using independently sampled noise trajectories and perturbed parameter values. The resulting temperature trajectories are used to construct full envelopes, showing the range within which the temperature may vary due to both internal stochastic variability and structural parameter uncertainty.
The parameters used in this simulation are: effective heat capacity $C=8\times 10^{8}\,\text{J/K/m}^{2}$, solar constant $S=1367.6\,\text{W/m}^{2}$, base albedo $\alpha=0.3$, and nominal emissivity $\epsilon=0.61$. The control input corresponds to an increase in albedo of $u(t)=0.2$, while the disturbance is modeled as a decrease in emissivity of $w(t)=-0.15$. The initial temperature is set to $T_{0}=288\,\text{K}$ (equivalent to $15^{\circ}$), and the simulation time span covers 30 years. Gaussian noise with a standard deviation of 10% is applied to the model dynamics, while the parameters $\alpha$ and $\epsilon$ are randomly perturbed with a standard deviation of approximately 1% around their nominal values to represent structural uncertainty. To maintain consistency with the rest of the tutorial, where temperature is presented in degrees Celsius, all plotted results are shifted from Kelvin to Celsius by subtracting 273.15. This makes the interpretation of the temperature trajectories more intuitive and aligned with conventional climate science practices.
It is important to emphasize that this example is conceptual in nature. The control input is not intended to represent a realistic or operational geoengineering scheme, nor does it account for complexities such as spatial heterogeneity, atmospheric chemistry, or long-term feedback mechanisms. Similarly, the temperature levels reached in the disturbed scenarios are unrealistically high relative to actual climate projections; they should not be interpreted literally but rather seen as illustrative. The objective here is not to simulate realistic climate outcomes, but to demonstrate how one might formally model, simulate, and reason about climate interventions using fundamental tools and concepts from control theory.
Having looked at the EBM from this perspective, some control-type questions can be posed such as: What input is needed to stabilize temperature at a target? How sensitive is the system to delays or uncertainty in the control action? These kinds of questions bring forward important concepts that connect climate modeling with insights from control engineering as discussed in Section VII.
V-C Interpreting Feedback in Climate Systems and Controls
One important concept here to discuss and be able to differentiate in climate modeling and control is feedback. In climate systems, feedback refers to how a change in one part of the system influences other components, which then either amplify or dampen the original change. A common example of positive feedback is the ice-albedo effect: warming leads to ice melt, which lowers Earth’s reflectivity (albedo), causing more solar absorption and further warming. On the other hand, negative feedback helps stabilize the system. For example, as temperature increases, the Earth emits more infrared radiation $\epsilon\sigma T(t)^{4}$, which counteracts the warming. In control engineering, feedback has a more formal definition: it refers to the process of measuring the output of a system and using that information to adjust the input. This can be either negative feedback, where the goal is to reduce the difference between a desired target and the actual output (stabilizing), or positive feedback, which reinforces changes and can lead to amplification or instability. In control terminology, feedback is part of the system architecture, often described using block diagrams or state-space representations, and is central to analyzing stability, robustness, and performance. What is referred to as feedback in climate modeling, such as ice-albedo or radiative cooling, corresponds in control terms to internal feedback loops in the plant or process model. These are natural feedback mechanisms built into the system’s dynamics, not part of a controller. Understanding this distinction is important when interpreting the behavior of the climate system from a control perspective—Fig. 8 presents a generalized diagram with the two feedback mechanisms illustrated.
Figure 8: Generalized feedback structure distinguishing internal and external feedback mechanisms. Internal feedback represents natural dynamics updated over time within the system—Feedback in climate science. External (human) feedback operates through measurement- and control-based approaches—Feedback in control engineering.
As a recap, we explain below how the dynamical nature of climate models connects to control concepts, using the EBM (1) as an example: The EBM is a nonlinear first-order dynamical system, where the state variable $T(t)$ evolves over time based on the net energy balance. The term $\frac{dT(t)}{dt}$ represents the time evolution of the system, which is central to both climate modeling and control theory. In its default form, the model is open-loop: it evolves based on physical parameters (e.g., solar radiation, albedo, emissivity) without any external regulation. Introducing a control input, such as stratospheric aerosol injection, allows us to influence the system—this effectively makes the model a closed-loop system. In climate models, feedbacks such as ice-albedo (positive) or radiative cooling (negative) are natural feedback mechanisms within the system—what control theory refers to as internal feedback in the plant dynamics. In control terminology, feedback typically refers to measuring the system output and adjusting the input accordingly, which can be done in climate policy by adapting actions based on observed temperature trends. Framing climate dynamics of this EBM in control terms allows us to ask questions like: What input is needed to stabilize temperature at a certain level? How does the system respond to delays or uncertainty in the input? What is the role of feedback in achieving or maintaining stability? This connection opens up new ways of analyzing and possibly influencing climate outcomes, using tools and insights from control engineering.
Beyond the simple EBM introduced earlier in this section, we present more complex models in the next section and discuss how they can be interpreted as control systems, along with the associated complexities. As examples, we focus on the dynamics of the atmospheric circulation, highlighting how this large-scale model can be viewed and analyzed within a control-oriented framework, providing mathematical formulations and discussing numerical solutions, feedback mechanisms, and parameterizations.
VI More Complex Dynamic Climate Model
As a more complex dynamic climate model, we consider a large-scale atmospheric motion model that is simulated using the primitive equations, derived from the Navier-Stokes equations under hydrostatic and shallow atmosphere assumption. This model captures the large-scale circulation processes in the atmosphere; see Tabs. I and II for relevance with other climate components and processes. The primitive model presented here captures the core dynamical structure of large-scale atmospheric motion. However, compared to comprehensive weather prediction systems, it neglects important moist processes, radiation, surface interactions, and chemical feedbacks. These additional processes, while increasing model complexity and computational cost, capture the full range of phenomena observed in weather systems—from convective storms to boundary-layer dynamics. These simplifications help isolate the fundamental atmospheric dynamical interactions between pressure, temperature, and wind, making the model particularly useful for conceptual understanding, mathematical analysis, and control-oriented formulations.
To that end, we do not present all the processes involved; rather, we describe the primitive model as a set of nonlinear PDEs that encode the conservation of momentum, energy, and mass. Then, a nonlinear ODE-based state-space representation is produced given the PDEs. The conservation laws are written in continuous form, assuming smooth spatial and temporal domains. The key state variables are the wind velocity $\mathbf{v}(\bm{r},t)$, temperature $T(\bm{r},t)$, and density $\rho(\bm{r},t)$, each modeled as a time-dependent field distributed over space, with spatial position denoted by $\bm{r}=[x,y,z]^{\top}$ and time by $t$. These quantities evolve dynamically through advection, external forcing, and internal physical interactions. Together, they form a compact state representation of the atmosphere’s large-scale dynamics, governed by the coupled system of equations shown as follows:
$\displaystyle\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}+f\mathbf{k}\times\mathbf{v}$ $\displaystyle=-\frac{1}{\rho}\nabla p+\mathbf{F}_{\text{viscous}},\leavevmode\nobreak$ (4a)
$\displaystyle\frac{\partial T}{\partial t}+\mathbf{v}\cdot\nabla T$ $\displaystyle=Q_{T},\leavevmode\nobreak$ (4b)
$\displaystyle\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{v})$ $\displaystyle=0.\leavevmode\nobreak$ (4c)
The wind velocity ${\mathbf{v}=[u,v,w]^{\top}}$ is a three-dimensional vector, with components in the zonal, meridional, and vertical directions, respectively. The temperature $T$, pressure $p$, and density $\rho$ are scalars. The Coriolis parameter is given by $f$, which depends on the Earth’s rotation rate and the latitude. The unit vector $\mathbf{k}$ points in the vertical direction. The term $\mathbf{F}_{\text{viscous}}$ represents viscous forces, and $Q_{T}$ denotes sources or sinks of temperature and heat (e.g., radiative or latent heating).
As mentioned, each field of states in this system is a function of space and time. Specifically, $\mathbf{v}(\bm{r},t)=\big{[}u(\bm{r},t),v(\bm{r},t),w(\bm{r},t)\big{]}^{\top}$, $T=T(\bm{r},t)$, $p=p(\bm{r},t)$, and $\rho=\rho(\bm{r},t)$, but we omit this explicit dependence in the notation for clarity. Similarly, the Coriolis parameter $f=f(\bm{r})$, which typically varies with latitude but may be assumed constant in simplified settings (e.g., the $f$-plane approximation). The unit vector $\mathbf{k}=[0,0,1]^{\top}$ points in the vertical direction and is constant in space and time. The terms $\mathbf{F}_{\text{viscous}}=\mathbf{F}_{\text{viscous}}(\bm{r},t)$ and $Q_{T}=Q_{T}(\bm{r},t)$ can vary in both space and time depending on the processes being modeled and approximations being adopted.
[…]
VI-A Understanding the Dynamics Behind the Atmospheric Primitive Equations
Having introduced the PDE (4), here we break down these systems of equations and how they interact with each other:
  1. 1.
    Momentum Equation (4a): governs the evolution of wind velocity $\mathbf{v}$ due to advection (nonlinear self-interaction of the wind velocity field $(\mathbf{v}\cdot\nabla)\mathbf{v}$), and the coriolis term $f\mathbf{k}\times\mathbf{v}$ accounts for Earth’s rotation. The pressure gradient force $\frac{1}{\rho}\nabla p$ drives wind from high to low pressure and the viscous terms $\mathbf{F}_{\text{viscous}}$ depict sub-grid friction or diffusion. The momentum equation is affected by density ${\rho}$ and pressure ${p}$, which are in turn influenced by temperature ${T}$ via the equation of state (ideal gas law) $p=\rho RT$, where $R$ is the specific gas constant.
  2. 2.
    Thermodynamic (Energy) Equation (4b): describes how temperature evolves due to (i) advection by wind $\mathbf{v}\cdot\nabla T$ (temperature is transported spatially) and (ii) heating term $Q_{T}$ which includes radiative heating, latent heat release, and sensible heat flux. Similarly, $\rho$ is affected via the ideal gas law, thus coupling back into the momentum equation.
  3. 3.
    Continuity Equation (4c): enforces mass conservation, links the density field to divergence/convergence in the flow ($\nabla\cdot(\rho\mathbf{v})$), and demonstrates how the changes in $\rho$ impact pressure gradients and therefore motion.
All three equations are nonlinearly coupled, meaning that a change in one variable affects the others. Additionally, uncertainties are inherently included in the parameters and unresolved terms. This is what gives atmospheric models their complex, often chaotic behavior. Notice that, these PDEs are based on the following assumptions:
  1. 1.
    Hydrostatic Approximation:
    • Assumes vertical pressure gradients balance gravitational acceleration, which is justified for large-scale motions where vertical accelerations are small (i.e., not deep convection): $\frac{\partial p}{\partial z}=-\rho g.$
    • Removes the vertical momentum equation from the full Navier-Stokes system.
    • The two previous assumptions greatly reduce numerical stiffness and computation time.
  2. 2.
    Shallow Atmosphere Approximation:
    • Assumes the thickness of the atmosphere is small compared to Earth’s radius.
    • Allows simplification in the geometry (e.g., assuming constant radius for Coriolis terms).
    • Horizontal metric terms (like curvature) are simplified.
VI-B Mathematical Interpretation of the Atmospheric Primitive Model
To facilitate interpretation for the reader and to support the later discussion on how they are solved, we break down each equation in (4) and explain the application of the relevant operators.
  1. 1.
    The Math in the Momentum Equation (4a):
    • $\displaystyle\dfrac{\partial\mathbf{v}}{\partial t}$ represents the dynamic temporal rate of change of the wind velocity field $\mathbf{v}$ at a fixed point in space. It is the time derivative of the state variable or the wind velocity.
    • $(\mathbf{v}\cdot\nabla)\mathbf{v}$ is the advection (nonlinear transport) term, which captures how a fluid parcel’s velocity changes as it moves through a spatially varying velocity field. Expanded component-wise as:
      $(\mathbf{v}\cdot\nabla)\mathbf{v}=\begin{bmatrix}u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}+w\frac{\partial u}{\partial z}\\ u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}+w\frac{\partial v}{\partial z}\\ u\frac{\partial w}{\partial x}+v\frac{\partial w}{\partial y}+w\frac{\partial w}{\partial z}\end{bmatrix}.$
      This term introduces nonlinear coupling among the velocity components.
[…]
    • $-\frac{1}{\rho}\nabla p$ is the pressure gradient force. A body force per unit mass that drives flow from high to low pressure regions. The gradient operator acts only on $p$ via $\nabla p=\begin{bmatrix}\frac{\partial p}{\partial x},\ \frac{\partial p}{\partial y},\ \frac{\partial p}{\partial z}\end{bmatrix}^{\top}$ and is scaled by the inverse of the local density $\rho$.
    • $\mathbf{F}_{\text{viscous}}$ is the viscous force term, representing internal friction in the fluid. It acts to diffuse momentum and slow down sharp velocity gradients. This force appears as a vector field with components in the $x$, $y$, and $z$ directions as follows $\mathbf{F}_{\text{viscous}}=\begin{bmatrix}F_{x},\ F_{y},\ F_{z}\end{bmatrix}^{\top}.$
    • Expanded Formulation: Putting all terms together, the momentum equation (4a) can be written in fully expanded matrix form as:
[…]
  2. 2.
    The Math in the Thermodynamic Equation (4b):
    • $\frac{\partial T}{\partial t}$ represents the local rate of change of temperature at a fixed point in space. It is the temporal derivative of the scalar state variable $T$.
    • $\mathbf{v}\cdot\nabla T$ is the advection of temperature by the velocity field. It is a directional derivative, expressing how temperature changes as a fluid parcel moves with the flow. Expanded as:
      $\mathbf{v}\cdot\nabla T=u\frac{\partial T}{\partial x}+v\frac{\partial T}{\partial y}+w\frac{\partial T}{\partial z}.$
      This term introduces coupling between the velocity field and the temperature field.
    • $Q_{T}$ represents external heating or cooling processes. This may include radiative heating, latent heat release, or sensible heat fluxes. It acts as a forcing term in the energy balance.
    • Expanded Formulation: Putting the terms together, the thermodynamic equation (4b) becomes:
      $\leavevmode\nobreak\ \frac{\partial T}{\partial t}=-\left(u\frac{\partial T}{\partial x}+v\frac{\partial T}{\partial y}+w\frac{\partial T}{\partial z}\right)+Q_{T}.$ (6)
  3. 3.
    The Math in the Continuity Equation (4c):
    • $\frac{\partial\rho}{\partial t}$ represents the local rate of change of density $\rho$ at a fixed point in space. This captures how density evolves over time due to fluid motion and compression or expansion.
    • $\nabla\cdot(\rho\mathbf{v})$ is the divergence of the mass flux. It accounts for the spatial variation in the flow of mass and can be expanded using the product rule as:
      $\nabla\cdot(\rho\mathbf{v})=\mathbf{v}\cdot\nabla\rho+\rho\nabla\cdot\mathbf{v},$
      where:
      $\nabla\cdot\mathbf{v}=\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}+\frac{\partial w}{\partial z}.$
      This term captures how variations in velocity and density contribute to local mass conservation.
    • Expanded Formulation: The continuity equation (4c) becomes:
[…]
Details on solving these equations and their discrete representation are provided in the following section.
VI-C Solving the Atmospheric Primitive Model
In summary, these equations simulate the evolution of wind velocity, temperature, and density fields in the atmosphere, governed by internal physical dynamics (advection, pressure gradients, and gravity). They also allow the incorporation of external forcings such as incoming solar radiation or anthropogenic GHGs concentrations as discussed later. Once formulated as a continuous set of those nonlinear PDEs [(5), (6), and (7)], the system can be discretized in space and time to be solvable numerically. We now discuss the spatio-temporal discretization resolution, initial and boundary conditions, integrating external forcing, parametrization, and feedback mechanisms in this model.
Tab. IV: Atmospheric Primitive Equation Terms and Operators: Units, Discrete Dimensions, and Physical Interpretation.
Symbol Units Discrete Dimension Represents Notes
$\mathbf{v}=[\mathbf{v}_{x},\mathbf{v}_{y},\mathbf{v}_{z}]$ [m/s] $\mathbb{R}^{3N}$ Wind velocity vector field Zonal (along $x$-axis), meridional (along $y$-axis), and vertical (along $z$-axis) wind components, stacked across grid points
$\mathbf{T}$ [K] $\mathbb{R}^{N}$ Temperature field Scalar field on the 3D spatial grid. Evolves due to advection and heating
$\bm{\rho}$ [kg/m3] $\mathbb{R}^{N}$ Air density Scalar field. Varies with altitude, pressure, and temperature. Appears in continuity and momentum equations
$\mathbf{p}$ [Pa] $\mathbb{R}^{N}$ Pressure field Drives wind through horizontal and vertical pressure gradients
$\nabla$ [1/m] $\mathbb{R}^{3N\times N}$ Gradient operator Maps scalar fields $\rightarrow$ vector fields. Constructed via finite-difference or finite-volume stencils
[…]
$\frac{\partial}{\partial t}$ [1/s] Time derivative operator Used in time integration. Realized through explicit or implicit stepping schemes
$f\mathbf{k}\times\mathbf{v}$ [m/s2] $\mathbb{R}^{3N}$ Coriolis acceleration term Represents Earth’s rotation effects. $f$ is Coriolis parameter and varies with latitude. $\mathbf{k}$ is unit vector in vertical direction. Cross-product per grid point
$\nabla\mathbf{p}$ [Pa/m] = [N/m3] $\mathbb{R}^{3N}$ Pressure gradient Spatial derivatives of pressure; drive acceleration in the momentum equation.
${\bm{\rho}}^{-1}\nabla\mathbf{p}$ [m/s2] $\mathbb{R}^{3N}$ Pressure gradient force Acceleration due to pressure gradient; nonlinearly depends on density.
$\mathbf{F}_{\text{viscous}}$ [m/s2] $\mathbb{R}^{3N}$ Sub-grid viscous force Represents turbulent diffusion or vertical mixing. Often parameterized via Laplacians.
$(\mathbf{v}\cdot\nabla)\mathbf{v}$ [m/s2] $\mathbb{R}^{3N}$ Advective acceleration Nonlinear advection of momentum by the wind field. Each component convects all others.
$\mathbf{v}\cdot\nabla T$ [K/s] $\mathbb{R}^{N}$ Advective heat transport Represents temperature change due to wind-driven advection.
$\mathbf{Q}_{T}$ [K/s] $\mathbb{R}^{N}$ Net heating source term Includes radiative, latent, and sensible heating. Can include sub-grid parameterized sources.
$\nabla\cdot(\bm{\rho}\mathbf{v})$ [kg/(m3·s)] $\mathbb{R}^{N}$ Mass divergence Appears in the continuity equation; determines local change in air density.
Discretization and Resolution. The atmospheric primitive equations (4) describe the evolution of velocity, temperature, and density fields over a continuous spatiotemporal domain. To solve them numerically, the spatial domain is discretized over a 3D grid, commonly defined in latitude-longitude-height or pressure coordinates, and the temporal dimension is discretized using time-stepping schemes.
[…]
  • Finite-difference methods (FDM): approximate derivatives using neighboring grid values (common in legacy models such as NCAR’s CAM [142]).
  • Spectral methods: decompose variables into basis functions (e.g., spherical harmonics). These methods are efficient for global models.
  • Finite-volume methods (FVM): integrate fluxes across control volumes; more robust for mass/energy conservation and common in newer models (e.g., FV3 in NOAA’s models [143]).
After discretizing the dynamics over the space domain, each field becomes a finite-dimensional vector:
$\mathbf{v}\in\mathbb{R}^{3N},\quad\mathbf{T}\in\mathbb{R}^{N},\quad\bm{\rho}\in\mathbb{R}^{N},\quad\mathbf{p}\in\mathbb{R}^{N}.$
In the continuous setting, the velocity field is written as $\mathbf{v}=(u,v,w)^{\top}$, where each component represents the zonal, meridional, and vertical wind velocity. After spatial discretization, each of these components becomes a vector in $\mathbb{R}^{N}$ defined at every grid point. The full velocity field is thus represented as a stacked vector $\mathbf{v}=[\mathbf{v}_{x}^{\top},\mathbf{v}_{y}^{\top},\mathbf{v}_{z}^{\top}]^{\top}\in\mathbb{R}^{3N}$, where $\mathbf{v}_{x}$, $\mathbf{v}_{y}$, and $\mathbf{v}_{z}$ correspond to the discretized fields for $u$, $v$, and $w$, respectively. In addition, the differential operators become sparse matrices or stencil-based operators (e.g., $\nabla\in\mathbb{R}^{3N\times N}$, $\nabla\cdot\in\mathbb{R}^{N\times 3N}$). For completeness, Tab. IV summarizes the system variables, parameters, and operators along with their units, discrete dimensions, and physical interpretations.
After spatial discretization, the PDEs (4) are transformed into a system of ODEs in time (semi-discrete form):
$\displaystyle\dot{\mathbf{v}}$ $\displaystyle=-(\mathbf{v}\cdot\nabla)\mathbf{v}-f\mathbf{k}\times\mathbf{v}-\bm{\rho}^{-1}\nabla\mathbf{p}+\mathbf{F}_{\text{viscous}},$ (8a)
$\displaystyle\dot{\mathbf{T}}$ $\displaystyle=-\mathbf{v}\cdot\nabla\mathbf{T}+\mathbf{Q}_{T},$ (8b)
$\displaystyle\dot{\bm{\rho}}$ $\displaystyle=-\nabla\cdot(\bm{\rho}\mathbf{v}).$ (8c)
These equations describe the evolution of state vectors $(\mathbf{v},\mathbf{T},\bm{\rho})$ governed by nonlinear and coupled dynamics. The nonlinear terms such as $(\mathbf{v}\cdot\nabla)\mathbf{v}$ and $\nabla\cdot(\bm{\rho}\mathbf{v})$ are evaluated using stencil operations or matrix–vector products. Similarly to (5), (6), and (7) in the continuous domain, the semi-discrete form can be expanded to:
[…]
where $\odot$ denotes element-wise multiplication across the $N$ grid points and the term $\bm{\rho}^{-1}$ represents the pointwise inverse of the discretized density field. That is, if $\bm{\rho}=[\rho_{1},\rho_{2},\dots,\rho_{N}]^{\top}$, then $\bm{\rho}^{-1}=[1/\rho_{1},1/\rho_{2},\dots,1/\rho_{N}]^{\top}$, and the expression $\bm{\rho}^{-1}\odot\nabla\mathbf{p}$ is computed element-wise across the grid.
Tab. V: Typical Spatial and Temporal Resolution for Global Atmosphere and Weather Models Built Based on Systems such as (8).
Model Type Horizontal Resolution Vertical Layers Time Step Notes
Global GCM (coarse) $\sim$100–250 km $\sim$20–60 layers 10–60 min Suitable for long-term atmospheric climate projections. Lower computational demand.
Global GCM (fine) $\sim$25–50 km $\sim$100+ layers 1–10 min Higher spatial fidelity allows better representation of changes in parameters and external forcings.
RCM / Weather Model $\sim$1–25 km $\sim$40–100+ layers seconds-minutes High-resolution simulations for localized weather or regional climate. RCM requires external boundary conditions (from GCM or reanalysis).
Temporal Discretization (Fully Discrete Form). Once the spatial discretization yields a system of ODEs (as in Eq. (8)), temporal integration is applied to evolve the system forward in time. This results in a fully discrete representation of the system dynamics, or a set of nonlinear difference equations (NDE).
The choice of time-stepping scheme influences numerical stability, accuracy, and computational cost:
  • Explicit methods (e.g., forward Euler, Runge-Kutta) evaluate the right-hand side of the ODE using known values at time step $n$. These methods are simple and computationally cheap but subject to stability constraints that may require very small time steps.
  • Implicit methods (e.g., backward Euler, Crank-Nicolson) involve evaluating the right-hand side at the unknown future time $n+1$, which generally requires solving a system of equations at each time step. They are more stable, especially for stiff systems, but computationally more intensive.
Let $\Delta t$ denote the time step, and define the time-discretized state variables as
$\mathbf{v}^{n}:=\mathbf{v}(t_{n}),\quad\mathbf{T}^{n}:=\mathbf{T}(t_{n}),\quad\bm{\rho}^{n}:=\bm{\rho}(t_{n}),$
$\text{where }t_{n}=n\Delta t\text{ and }n=0,1,\dots,N_{t}.$
Meaning that, the simulation is carried out over a time span $[0,t_{\text{final}}]$, where $t_{\text{final}}=N_{t}\Delta t$, and $N_{t}$ is the total number of time steps.
Using a forward Euler scheme as an illustrative example (first-order explicit method), the time integration of Eq. (8) gives the fully discrete update equations:
$\displaystyle\mathbf{v}^{n+1}$ $\displaystyle=\mathbf{v}^{n}+\Delta t\cdot\Big{(}-(\mathbf{v}^{n}\cdot\nabla)\mathbf{v}^{n}-f\mathbf{k}\times\mathbf{v}^{n}$
$\displaystyle\quad-\bm{\rho}^{-1,n}\odot\nabla\mathbf{p}^{n}+\mathbf{F}_{\text{viscous}}^{n}\Big{)}$ (10a)
$\displaystyle\mathbf{T}^{n+1}$ $\displaystyle=\mathbf{T}^{n}+\Delta t\cdot\Big{(}-\mathbf{v}^{n}\cdot\nabla\mathbf{T}^{n}+\mathbf{Q}_{T}^{n}\Big{)}$ (10b)
$\displaystyle\bm{\rho}^{n+1}$ $\displaystyle=\bm{\rho}^{n}+\Delta t\cdot\Big{(}-\nabla\cdot(\bm{\rho}^{n}\mathbf{v}^{n})\Big{)}$ (10c)
where the superscript $n$ indicates the current time step, and $n+1$ the next; all nonlinear terms are evaluated at time step $n$, consistent with the explicit Euler scheme; the inverse density term $\bm{\rho}^{-1,n}$ is computed element-wise as in the semi-discrete form: if $\bm{\rho}^{n}=[\rho_{1}^{n},\dots,\rho_{N}^{n}]^{\top}$, then $\bm{\rho}^{-1,n}=[1/\rho_{1}^{n},\dots,1/\rho_{N}^{n}]^{\top}$. Finally, the pressure gradient $\nabla\mathbf{p}^{n}$ and all spatial derivatives are approximated using the same stencil or matrix-based discretization discussed earlier.
This formulation yields a first-order, fully discrete system that updates the state at each time step based solely on known quantities. Higher-order or implicit schemes can be applied in similar fashion, but may require iterative solvers or linearizations.
Discretized System’s Resolution. The spatial and temporal resolutions (e.g., the grid size and time step) are determined based on the type of model as shown in Tab. V and the discretization scheme used, with stability and accuracy constraints taken into account. The reader is also referred to Tab. III for more details on those climate models types. The choice of resolution reflects a tradeoff between computational tractability and the ability to represent fine-scale processes, which is also influenced by the model’s scientific objectives (e.g., climate trends vs. short-term forecasting) and the availability of computational resources. As such, the structure and interpretation of the discrete equations (10) presented here remain applicable across a wide range of model configurations, from coarse-resolution climate simulations to high-resolution, event-driven forecasts.
Initial vs. Boundary Conditions (Weather vs. Regional Climate Models). Global weather models obtain their forecasts by solving an atmospheric dynamical system (e.g., (10)) forward from a detailed initial atmospheric state (at $t=0$). On the other hand, regional weather/climate models (based on these same dynamics) solve them with boundary conditions applied at the edges of the domain: $\mathbf{v},\mathbf{T},\bm{\rho}$ prescribed at domain boundaries every 3–6 hours, to ensure consistency with large-scale atmospheric patterns. These boundary values are interpolated from a GCM or reanalysis dataset, and they feed into the right-hand sides of (10) via advection or pressure gradient terms.
Parameterization and External Forcing. In Eq. (10), the term $\mathbf{F}_{\text{viscous}}$ represents unresolved turbulent diffusion and frictional effects. Since many atmospheric motions (e.g., small eddies and turbulence) occur at spatial and temporal scales smaller than the grid resolution, they cannot be explicitly modeled and must be parameterized. Therefore, it is counted as an unresolved processes and its effect is approximated based on a resolved state. For example, turbulent mixing in the planetary boundary layer is often modeled using eddy diffusivity schemes, which in return can be expressed as function of the wind velocity $\mathbf{v}(t)$.
External forcing refers to influences and external inputs to that affect the atmospheric state but are not determined by the internal dynamics of the system. Common examples for the systems expressed in (4) and (10) include solar radiation (shortwave), GHGs concentrations, aerosols, land use changes, and sea surface temperature boundary conditions. These forcings typically enter the system through the heating term ($Q_{T}$ in (4) and $\mathbf{Q}_{T}$ (10)) and/or through modified boundary conditions for regional models.
Feedback Mechanisms. Even in this core dynamic system, internal feedback mechanisms naturally emerge from the coupling of variables:
  • Thermodynamic–momentum coupling: A change in temperature $T$ modifies air density $\rho$ via the ideal gas law, which in turn alters pressure gradients $\nabla p$ and affects wind velocity $\mathbf{v}$. The wind field then advects temperature, feeding it back into the temperature field and completing the loop as shown below:
    $T$$\rho$$\nabla p$$\mathbf{v}$$\nabla T$$T$
  • A heating anomaly ($Q_{T}>0$) increases the air temperature $T$, which reduces density $\rho$ through the ideal gas law and alters pressure gradients $\nabla p$. This accelerates the wind field $\mathbf{v}$, which transports heat and moisture, further modifying the temperature field. The increase in $T$ also raises water vapor concentration, a GHG, enhancing greenhouse trapping and amplifying the initial warming via radiative feedback. These couplings form a closed internal feedback loop in the atmospheric system:
    $Q_{T}>0\rightarrow T\rightarrow\rho\rightarrow\nabla p\rightarrow\mathbf{v}\rightarrow\nabla T\rightarrow$ GHG forcing $\rightarrow Q_{T}$
VI-D State-Space Outlook on the Atmospheric Model
Having introduced the structure, physical, and mathematical interpretation of the atmospheric primitive equations, we now shift toward viewing this system through the lens of control engineering. While originally formulated to simulate large-scale atmospheric dynamics, this model can also be understood as a high-dimensional, nonlinear dynamical system with coupled state variables, external inputs, and internal feedback loops. By identifying the system’s states, inputs, outputs, and feedback structure, we lay the groundwork for applying tools from control engineering such as observability, controllability, stability, state estimation, and feedback control design. This enables potentially improved understanding, analysis, and prediction of the behavior of the atmospheric system.
The handling of the atmospheric primitive system pipeline progresses through three stages:
[…]
where $\mathbf{x}$ is the full state vector gathering the wind velocity $\mathbf{v},\mathbf{T},$ and $\bm{\rho}$.
We can now reinterpret the atmospheric model using a state-space framework commonly adopted in control engineering. This provides a compact and general way to represent system dynamics, inputs, outputs, and parameters.
State-Space Form of the Semi-Discrete ODE System. The system of equations in (8) describes the evolution of the atmospheric state variables over time following spatial discretization. This system can be expressed in a general nonlinear state-space form:
$\displaystyle\frac{d\mathbf{x}}{dt}=\mathbf{f}(\mathbf{x}(t),\,{\mathbf{u}(t)},\,{\mathbf{w}(t)},\,{\bm{\theta}(t)};\,\mathbf{c}),$ (11a)
$\displaystyle\mathbf{y}(t)=\mathbf{h}(\mathbf{x}(t),\,\mathbf{u}(t),\,\mathbf{w}(t),\,\bm{\theta}(t)).$ (11b)
In this representation, $\mathbf{x}(t)=[\mathbf{v}_{x}(t)^{\top},\,\mathbf{v}_{y}(t)^{\top},\,\mathbf{v}_{z}(t)^{\top},\,\mathbf{T}(t)^{\top},\,\bm{\rho}(t)^{\top}]^{\top}\in\mathbb{R}^{5N}$ is the state vector, composed of the time-varying discretized fields of wind, temperature, and density. The input $\mathbf{u}(t)$ denotes potential control actions, such as prescribed heating, anthropogenic forcing, or modified boundary fluxes. The term $\mathbf{w}(t)$ represents uncontrolled environmental inputs like volcanic eruptions or natural solar variability. The time-dependent parameter vector $\bm{\theta}(t)$ includes evolving but externally provided fields such as surface albedo, land cover, or radiative coefficients. Lastly, $\mathbf{c}$ represents time-invariant physical parameters (e.g., gas constant, gravity), and the output vector $\mathbf{y}(t)$ captures observable quantities such as surface temperature, pressure, or wind components at selected locations or aggregated regions.
Fully Discrete Nonlinear Dynamical System. By further discretizing the temporal domain using a numerical integration scheme such as forward Euler (as in (10)), the system becomes a NDEs:
$\mathbf{x}^{n+1}=\mathbf{f}_{d}(\mathbf{x}^{n},\,\mathbf{u}^{n},\,\mathbf{w}^{n},\,\mathbf{p}^{n};\,c)$ (12)
$\mathbf{y}^{n}=\mathbf{h}_{d}(\mathbf{x}^{n},\,\mathbf{u}^{n},\,\mathbf{w}^{n},\,\mathbf{p}^{n})$ (13)
where $\mathbf{x}^{n}\in\mathbb{R}^{5N}$ is the state at time $t_{n}=n\Delta t$, all variables are evaluated at discrete time steps indexed by $n$, and the function $\mathbf{f}_{d}$ results from applying the chosen time-stepping scheme to (11).
Having reformulated the atmospheric dynamics in both continuous and discrete state-space form, we are now equipped to draw explicit parallels with the standard language and structure of control systems. This interpretation is not only useful for re-expressing the model mathematically, but it also enables the application of a wide array of techniques from estimation, feedback control, and system analysis. The following summary boxes highlight how the atmospheric primitive equations, originally derived for physical simulation, can be naturally framed as a large-scale nonlinear dynamical system amenable to control-theoretic reasoning.
[…]
In GCMs, which typically operate with horizontal resolutions between 100 and 250 kilometers (or down to 25-50 km in higher-resolution configurations as shown in Tab. V), the total number of state variables can range from tens to hundreds of millions, depending on the number of vertical levels and included fields. For example, a typical GCM might use a horizontal resolution of 100 km with 60 vertical levels on a grid covering the entire globe ($3.6\times 10^{6}$ horizontal grid points). Tracking three primary fields of states: wind velocity $(\mathbf{v}_{x},\mathbf{v}_{y},\mathbf{v}_{z})$, temperature $\mathbf{T}$, and density $\bm{\rho}$ at each of the  $2.2\times 10^{8}$ total grid points (horizontal $\times$ vertical) results in over one billion state variables.
RCMs and NWP systems, while applied over smaller domains, use much finer resolution, often between 1 and 25 kilometers, and thus also produce systems with millions of coupled state variables. Another example here is a RCM applied over a smaller domain (e.g., $1000\times 1000$ km) at 10 km resolution with 50 vertical levels would yield $10^{6}$ horizontal points and about $5\times 10^{7}$ total variables for the same set of fields. Reducing the horizontal resolution to 5 km would quadruple the number of horizontal points and increase the system size accordingly.
These discretized systems retain the nonlinear and tightly coupled structure of the original equations, making them ideal candidates for control-theoretic interpretation and analysis, especially in the context of large-scale estimation, uncertainty quantification, and feedback structure.
[…]
Raising Control Engineering Questions. The models presented in (4) and (10) open up classical and modern control questions that are related to core control concepts, such as the following, pertaining to specifically observability and controllability.
  • Observability and Sensor Placement: Which parts of the atmospheric state can be reconstructed from the available measurements? For specific kind of sensors, what are the unobservable subspaces in one specific climate model? What and how many sensors are required to ensure desirable observability metrics for any of the derived climate models? These questions are often posed in dynamic system sciences but are hardly navigated in climate research, in this particular system-theoretic context.
  • Controllability, Reachability Analysis, and Control Node Placement: If hypothetical actuators are introduced to imitate controlled practices (e.g., geoengineering techniques) and policy-driven mitigation schemes, then mapped to present control vectors in (4) and (10), then how effective are these control measures at steering the system toward a more desirable climate trajectory? What are the resulting controllable subspaces and how can reachability analysis improve our understanding of the decade-long evolution of climate states? Instead of just simply running extended period simulations of climate models, can reachability analysis inform or predict reachable spaces given improved mitigation strategies and specific control actions? What is the impact of placing control nodes as actuators in specific geographic locations to thereby improve the trajectory of climate states? Can this be posed as an actuator placement problem?
While these questions are intuitive to the controls community, they may appear unfamiliar or even abstract to climate scientists. However, many of these concepts have meaningful analogues in climate modeling, even if they are approached from different directions or embedded implicitly in other tasks. This connection is further elaborated and emphasized in Section VII.
VII Analogies Between Control Engineering and Climate Science Problems
In this section, we draw and observe structured connections between challenges (and research problems) in climate science and foundational ideas in control engineering. The goal is to suggest that some known climate problems can be potentially cast as control engineering ones. By using the conceptual lexicon of control engineering (states, inputs, outputs, observability, controllability, stability feedback, model order reduction, system identification, and uncertainty propagation), we build a shared framework for interdisciplinary thinking. We argue here that control systems researchers can benefit from the produced analogies in this section by hopefully applying some control engineering algorithms for large-scale climate models. We admit that this might be very ambitious, but the point of this tutorial was to initiate a conversation and hopefully motivate some folks to delve into climate science and seek collaborators from this crucial field of science.
The previous sections have laid the groundwork for this discussion by emphasizing the dynamic and time-evolving nature of climate models. We have introduced a simple EBM, a first-order nonlinear dynamical system representing the evolution of global mean temperature. This model serves as a starting point to illustrate how climate behavior can be framed in terms of input-output relationships, feedback mechanisms, and externally driven disturbances. We have introduced hypothetical control inputs, such as geoengineering strategies to increase planetary albedo, and external disturbances, i.e., radiative forcing from greenhouse gas emissions, to demonstrate how climate models can be extended into open- or closed-loop control systems. These formulations naturally raise control-oriented questions regarding stabilization, responsiveness, sensitivity, and robustness, even in their simplest form.
To go beyond the conceptual ODE-level intuition, we then have introduced a more complex and physically realistic climate model: a large-scale atmospheric system based on the primitive equations. These nonlinear PDEs govern the evolution of wind, temperature, and density fields under the conservation of momentum, mass, and energy. After spatial and temporal discretization, the resulting system becomes a high-dimensional nonlinear dynamical model with millions of coupled variables. This transformation highlights the rich dynamical structure of climate systems and presents a direct opportunity to apply control-theoretic reasoning. The primitive model exposes the system’s internal feedbacks, distributed inputs and outputs, external forcing, and sensitivity to uncertainty—key features that align closely with control formulations.
Building on these conceptual extensions, this section formalizes and expands the parallels between climate modeling and control system analysis. Control engineering provides a rich set of tools for modeling, analyzing, and designing dynamical systems under uncertainty, partial observability, and feedback. These are exactly the features that characterize large-scale climate systems. While traditional climate models are built from first principles, they often face challenges that align closely with those addressed by control theory: high dimensionality, sparse data, parameter uncertainty, delayed responses, and the need for optimal intervention under constraints.
Rather than treating climate models as black boxes or fixed simulation tools, the control perspective views them as dynamical systems whose behavior can be shaped, predicted, and interpreted through formal system-theoretic properties. In doing so, it becomes possible to translate familiar engineering tasks, such as state estimation, feedback design, reachability, or robustness analysis, into meaningful questions within climate science. We do not imply here that control theory will change the discourse of climate science; we argue that control engineers and researchers can perhaps learn from climate problems and perhaps be more involved in the most timely research topic of our time.
This section presents a structured exploration of these connections, with the aim of fostering cross-disciplinary insight and enabling the use of control engineering frameworks to deepen our understanding of climate dynamics and intervention planning. That is, the remainder of this section highlights several core concepts in control theory and their relevance to climate modeling. Tab. VI summarizes these parallels.
Tab. VI: Parallels Between Climate Science and Control Engineering, Emphasizing Shared Challenges and Transferable Tools Across Disciplines.
Climate ScienceControl EngineeringProblemDefinitionProblemDefinition Missing Data & Data Assimilation Data assimilation merges model forecasts with sparse or noisy observations to estimate the full system state. This enables better initialization of forecasts, especially when key variables are unobserved or partially measured. Observability & State Estimation Control systems use observer design and estimation filters (e.g., Kalman filters) to reconstruct internal states from outputs. These techniques parallel climate data assimilation, offering formal tools for state reconstruction and uncertainty handling. Climate Intervention Planning Climate interventions aim to alter system trajectories, e.g., reducing global temperature or emissions. The effectiveness of these actions depends on how responsive the climate system is to applied forcing, much like input response in control systems. Controllability of
Dynamical Systems
Controllability characterizes whether and how a system can be steered to desired states using inputs. Tools such as input Gramians, metrics, and actuator design offer formal ways to plan effective interventions: concepts increasingly relevant in climate control strategies. Policy Levers &
Climate Goals
Climate actions like carbon pricing, energy transition, or geoengineering serve as levers to steer the system toward goals (e.g., net-zero emissions, temperature caps). Planning these involves choosing when and how intensely to act, under uncertainty and constraints. Optimal Control &
Trajectory Design
Optimal control computes input trajectories to meet system objectives while respecting dynamics and constraints. This maps to climate policy design, where actions are optimized over time to achieve climate targets under physical and socioeconomic limits. Forecasting &
Climate Projections
Forecasts estimate short-term evolution (e.g., seasonal rain patterns), while projections map long-term plausible futures under various scenarios. Both explore how the system might evolve from today’s conditions under different influences. Reachability Analysis Reachability defines the set of states a system can reach given initial conditions and inputs. Like climate projections, it maps possible futures in forms of envelopes rather than certainties, helping guide planning under uncertainty. Uncertainty Quantification
& Climate Risk
Uncertainty is central to climate modeling, arising from internal variability, incomplete data, and model limitations. Climate scientists use ensemble projections, confidence intervals, and scenario spreads to assess how robustly future states can be anticipated. Robustness &
Uncertainty Handling
In control engineering, robustness ensures that systems perform reliably under uncertainty in dynamics or disturbances. Techniques like $H_{\infty}$ control, tube MPC, and bounded-error design echo ensemble-based and worst-case strategies in climate risk management. Tipping Points
and System Resilience
Climate systems may experience abrupt, irreversible shifts if critical thresholds are crossed. Rather than returning to a pre-disturbance state, resilience analysis focuses on maintaining the system within safe functional regimes and avoiding destabilizing feedback loops. Stability and
Dynamical Behavior
In control theory, stability ensures that a system returns to equilibrium after perturbations. Techniques such as Lyapunov analysis, eigenvalue tests, and bifurcation theory characterize how systems behave near equilibria, concepts that help frame climate tipping dynamics and resilience thresholds. Model Calibration
& Learning
Model calibration tunes parameters of dynamical models (e.g., cloud, land, ocean processes) to better fit observations. Learning techniques like ML and ensemble methods have been used to infer subgrid processes, design surrogate models, and replace or enhance traditional physical parameterizations, while aiming to maintain physical consistency. System Identification
& Model Discovery
System identification builds mathematical models from input-output data. Techniques range from ARX models and subspace identification for linear systems to modern approaches like sparse identification (SINDy) and physics-informed ML for nonlinear systems. Hybrid modeling blends physical knowledge with learned components to improve prediction and control.
VII-A Missing Data Points & Data Assimilation vs. Observability & State Estimation
We start out the technical details in this section by building parallels between the climate problems of missing data points and data assimilation versus the control problems of observability and state estimation.
The challenge of reconstructing the system parameters and state from partial or missing data arises in both climate science and control engineering, though each field has developed its own terminology and tools. In control engineering, observability refers to the ability to infer the internal state of a dynamical system using only output measurements. Building on this, state estimation techniques, such as Kalman filtering or observer design, are used to reconstruct those internal states over time, even in the presence of noise or data loss [144, 145, 146, 147, 148, 149, 150, 151, 152]. In climate science, similar problems are addressed under the umbrella of data assimilation [153, 154, 155, 156, 157, 158, 159, 160]. Here, incomplete or uncertain observational datasets (e.g., from satellites or ground sensors) are merged with model predictions to yield improved estimates of climate variables such as temperature, pressure, or humidity. This is especially critical when direct observations are sparse or missing over space or time. The parallels between these two domains enable a cross-disciplinary perspective. We produce particular examples and discussions.
VII-A1 Kalman Filtering and Data Assimilation
The Kalman filter, a foundational tool in control engineering for linear Gaussian systems, has been adapted in climate science as the basis for various data assimilation schemes [161, 162, 163, 164, 165, 166, 167]. Ensemble Kalman Filters and extended Kalman filters are direct generalizations that handle nonlinear climate models and ensemble-based uncertainty quantification. An example of this adaptation is the assimilation system developed by Koshin et al. [162], where a four-dimensional local ensemble transform Kalman filter (4D-LETKF) is used with a high-top general circulation model to reconstruct the atmospheric state up to the lower thermosphere. Observations assimilated include conventional ground-based measurements from available datasets and satellite temperature profiles from the Aura Microwave Limb Sounder (MLS), covering heights from the troposphere through the mesosphere. This has enabled state estimation across scales, from synoptic dynamics to mesospheric processes, critical for capturing phenomena such as sudden stratospheric warmings (SSWs) and inter-hemispheric coupling.
Another example is the work by Pedatella et al. [165], who have applied an ensemble adjustment Kalman filter (EAKF) within the whole atmosphere community climate model (WACCM) framework. Their system assimilates satellite observations to improve simulations of large-scale events like the 2009 sudden stratospheric warming. Observations include satellite temperature and ozone data, and the model has spanned the entire atmosphere up to 140 km altitude, demonstrating how Kalman-based assimilation could improve the consistency and predictability of climate variables even under extreme dynamical disturbances.
The utilization of basic Kalman filtering techniques illustrates that there is an opportunity for control engineers to: (i) test out advanced filtering and state estimation algorithms (such as advanced observers, moving horizon estimators, learning-based predictors/estimations for large-scale nonlinear climate models); (ii) develop new state estimation algorithms that perform well for climate models with real data. This is because it indeed is known that the basic Kalman filter—and its many variants—suffer in handling unmodeled uncertainty, complex nonlinearities, and unknown initial conditions.
VII-A2 Handling Missing Data
In control theory, missing or unmeasured outputs are often managed through robust estimation algorithms and strategic sensor placement to maximize system observability. Observers are carefully designed to compensate for the lack of complete output information, ensuring reliable state recovery even with incomplete sensing. In climate science, missing observations can arise due to sensor malfunctions, satellite coverage gaps, or inaccessible regions (e.g., polar areas, deep oceans). These gaps are typically addressed by interpolation, reanalysis products, or machine-learning-based gap-filling techniques. As an example, reanalysis datasets like ERA5 [168] merge sparse, noisy measurements with physical models to reconstruct missing atmospheric and oceanic fields over decades. The missing data problems in climate science are important because they allow: (i) reconstructing or estimating past climate trajectories, (ii) validating physics-based climate models (such as the ones outlined earlier in the paper), and (iii) accurately predicting with some certainty future climate. All of these climate problems and outcomes (i)(iii) are common in control systems expositions and hence pursuing them in a control lens can be fundamentally new—and potentially useful.
VII-A3 Observer Design and Climate States Reconstruction
Observer theory provides formal designs for estimating the hidden state of a system using known inputs and outputs. Analogously, climate models can integrate these principles to estimate unobserved fields (e.g., subsurface ocean temperature) using observable proxies (e.g., sea surface temperature), thereby improving model fidelity and forecast accuracy.
These connections observed in this section underscore a key opportunity: methods from control engineering, grounded in estimation theory, observability analysis, and feedback structure, can be tailored and extended to support climate science needs in data-limited regimes. By bridging these frameworks, we enable more robust and theoretically sound strategies for climate parameters and state reconstruction and prediction.
VII-B Climate Interventions, Policy Levers, & Climate Goals vs. Controllability & Optimal Control
Having thoroughly explained connections between state estimation and data assimilation in climate science, it is only natural to pose questions pertaining to controllability analysis and its corresponding relevance in climate studies. The question of how much influence we can exert on a system, whether to guide it toward a desired outcome or away from a harmful trajectory, is a shared challenge in both climate science and control engineering. While control theory provides formal tools to assess this influence through controllability analysis, climate science deals with it more implicitly, through studies on the effectiveness and consequences of human interventions such as emissions reductions or geoengineering.
In control engineering, controllability refers to the ability to drive the system from any initial state to any final state using admissible inputs. This is analyzed using mathematical tools such as the controllability matrix, Gramian analysis, metrics, and actuator placement strategies. Inputs are well-defined and can be applied selectively in space and time to shift the system in desired directions [145, 144, 148]. In climate science, interventions represent policy or engineering efforts to alter the trajectory of climate variables—for example, by reducing CO2 emissions, increasing surface albedo, or deploying stratospheric aerosols [169, 170, 171]. These are analogous to control inputs, as introduced in Section V and reflected by introducing aerosol injections as control input in the simple EBM model (2). However, unlike classical control systems, the input-output pathways in climate models are uncertain, delayed, and often spatially heterogeneous.
Despite these differences, several concepts in control engineering can inform and enhance how we model, evaluate, and plan climate interventions. Framing climate action as a control problem allows us to systematically assess not just what interventions are possible, but how effective and efficient they are likely to be, and where control energy may be wasted.
Figure 9: Projected net global GHG emissions under various policy scenarios. The red shading shows expected GHG emissions based on currently implemented policies, while the green and blue pathways represent the more aggressive reductions needed to limit global warming to 2C and 1.5C, respectively. Shaded regions represent uncertainties in model projections (5–95% confidence intervals). The data illustrate how different levels of intervention can significantly alter emissions trajectories. Figure adapted from the IPCC Sixth Assessment Report (AR6) [15]. The figure sheds light into the potential role of control in climate science. Current mitigation trajectories are not designed through formal optimal control formulations where climate dynamics are used directly to compute minimum-control strategies. In contrast, optimal control theory offers a systematic framework for designing interventions: defining explicit objectives, modeling system dynamics, and optimizing trajectories. By framing climate action through the lens of control, we can move beyond ad-hoc aggregate targets and toward deeply planned, dynamically feasible pathways that respect both physical and socioeconomic constraints. A pertinent investigation is hence whether plots such as this one can be generated via specific control actions, and/or ranges for states can be computed via reachability analysis considering a fixed set of mitigation strategies and control signals.
As an illustrative example, Fig. 9 shows another important dimension: the projected pathways for global GHG emissions under different levels of policy action. Current policies, as shown by the red band, are not sufficient to meet the targets associated with limiting warming to 1.5C or 2C. Achieving those targets would require far more aggressive and coordinated emissions reductions, as depicted by the green and blue trajectories. Importantly, these emissions scenarios are typically constructed and simulated by aggregating policy pledges and expert assumptions, rather than by solving climate dynamics-constrained optimization or optimal control problems. That is, current mitigation trajectories are not designed through formal optimal control formulations where climate dynamics (governed by ODEs/PDEs such as the models presented in Sections V and VI) are used directly to compute minimum-control strategies to stabilize the system at target states (e.g., CO2 concentration or surface temperature). In contrast, optimal control theory offers a systematic framework for designing interventions: defining explicit objectives (e.g., limiting temperature rise), modeling system dynamics, applying control inputs (e.g., emissions reductions), and optimizing trajectories over time while satisfying system constraints. This perspective not only enables finding feasible mitigation pathways but also identifies trade-offs, energy costs, and the sensitivity of outcomes to uncertainties. By framing climate action through the lens of dynamic systems and control theory, we can move beyond ad-hoc aggregate targets and toward deeply planned, dynamically feasible pathways that respect both physical and socioeconomic constraints.
That is, the challenge in climate science involving determining what actions should be taken, when, and at what intensity is a structure that directly parallels optimal control problems in engineering. In both domains, the objective is to influence the evolution of a dynamical system in order to achieve certain outcomes, such as minimizing a cost or reaching a target state.
In control engineering, optimal control provides a framework to compute input trajectories that minimize an objective function while satisfying system dynamics and constraints. Classical examples include minimizing energy use, time to reach a target, or deviation from a reference trajectory. In climate science, this maps to the design of long-term policy pathways, where actions such as emissions reduction, energy transition, or geoengineering are applied over time to meet targets like temperature stabilization or CO2 budgets. These policies act as control inputs, while climate models represent the system dynamics.
Several concepts in optimal control have natural analogues in climate decision-making:
  • Objective Functions: Control systems minimize energy, cost, or error. Climate policy optimizes for carbon budgets, economic cost, risk, or social impact which often use IAMs—refer to Tab. III.
  • Constraints: Just as control problems enforce bounds on states and inputs, climate policies must respect physical, political, and technological limits (e.g., renewable deployment rates, economic costs, fairness constraints).
  • Control Horizon and Receding Horizon Control: Optimal control often looks ahead over a fixed time window. Climate policy planning does the same, with tools like model predictive control (MPC) increasingly adapted to accommodate feedback, uncertainty, and mid-course correction.
  • Trade-offs and Multi-Objective Control: In both domains, competing goals (e.g., economy vs. environment, short-term vs. long-term) are common. Analyses such as Pareto front and weighted cost formulations help frame these trade-offs mathematically.
Other problems and analogies are still relevant.
VII-C Forecasting & Prediction vs. Reachability Analysis
Forecasting in climate science and reachability analysis in control engineering address a common question: What future states can the system reach from its current condition? In climate science, a distinction is made between forecasts and projections [172, 173, 174, 175]. Forecasts are short- to medium-term weather predictions (typically days to seasons) of climate variables like temperature or precipitation, based on current initial conditions and dynamically evolving physical models, much like state prediction in control systems. These forecasts aim to be precise and are evaluated against reality. In contrast, projections refer to long-term simulations that estimate how the climate system might evolve under different assumed external forcings or socio-economic pathways, such as varying greenhouse gas emission scenarios. These are often simulated as not direct predictions but conditional if-then explorations of plausible futures.
In control engineering, reachability analysis plays a parallel role by determining all states that a system can attain over time given an initial condition and a set of admissible inputs [176, 177, 178]. It provides bounds on what is possible, rather than what is certain; much like how climate projections outline the range of possible future climates under specific pathways. The ensemble of climate model outputs from scenarios like those in CMIP6 (e.g., Shared Socio-economic Pathways SSP1-2.6, SSP5-8.5) creates a projection envelope that is conceptually similar to a reachable set in control theory. Both approaches aim to map out a future space that informs planning and decision-making under uncertainty. Importantly, while traditional climate projections require computationally intensive ensemble simulations across numerous scenarios, reachability analysis offers a promising alternative. By characterizing the entire set of possible future states using system dynamics and input bounds, it allows us to infer whether critical or dangerous states, such as crossing a temperature threshold or triggering a tipping point, could be reached. This approach can potentially, in principle, circumvent the need to re-run large-scale models for every new scenario, providing guarantees over the space of uncertainty without exhaustive simulation. To that end, key parallels between the domains include:
  • Both rely on system dynamics and are sensitive to initial state uncertainty and input variability.
  • Reachable sets in control match the role of scenario-based projections in climate modeling.
  • Techniques used in MPC for safety-constrained forecasting have conceptual overlap with constrained climate mitigation planning.
These parallels can be further pursued and rigorously investigated.
VII-D Uncertainty Quantification & Climate Risk vs. Robustness & Decision Confidence
Uncertainty is a central theme in both climate science and control engineering. In each field, the presence of imperfect knowledge, noisy observations, and model approximations demands tools and frameworks that explicitly account for uncertainty in both analysis and decision-making. Although the terminology and contexts differ—climate scientists speak of ensemble projections and confidence intervals, while control engineers use concepts like robustness and bounded uncertainty—the underlying goal is the same: to understand the limits of what can be known and to make decisions that remain valid across a range of possible conditions.
In control engineering, robustness refers to the ability of a system or controller to maintain acceptable performance despite uncertainties in model parameters, external disturbances, or sensor noise. This includes classical approaches like $\mathcal{H}_{\infty}$ control, min-max optimization, and tube-based MPC, all of which aim to ensure reliable operation under worst-case or bounded uncertainty [179, 180, 181, 182]. Meanwhile, climate science approaches uncertainty through formal uncertainty quantification (UQ), using statistical, stochastical, and ensemble-based methods to capture the range of possible future outcomes. This includes multi-model ensembles (e.g., CMIP6), Monte Carlo sampling, and Bayesian frameworks to characterize and communicate confidence in projections [183, 184, 185, 186, 187].
Despite disciplinary differences, the treatment of uncertainty in both fields reveals strong structural parallels:
  • Uncertainty Sets and Ensembles: In robust control, uncertainty is represented by structured sets (e.g., intervals, polytopes, ellipsoids). In climate modeling, ensembles of models or parameter sets perform a similar role, each representing a plausible system realization.
  • Disturbance Rejection vs. Scenario Spread: Robust controllers are designed to reject or attenuate unknown inputs. Climate projection ensembles under varying scenarios (e.g., SSP pathways) represent a spread in potential external forcings, highlighting sensitivity to input uncertainty.
  • Safety Margins and Exceedance Thresholds: Control systems include margins to ensure safety under worst-case disturbances. Similarly, climate policy aims to avoid crossing critical thresholds (2°C warming), expressed in terms of exceedance probabilities or confidence bounds.
  • Decision-Making under Uncertainty: Both fields rely on optimization techniques that explicitly consider uncertainty, such as stochastic MPC in control or robust emissions pathways in climate IAMs.
By recognizing uncertainty as a shared, intrinsic feature of dynamical systems, we can better appreciate the synergy between climate modeling and control theory. Tools from both domains can inform one another, not through transplantation but through analogy.
VII-E Tipping Points & System Resilience vs. Stability & Dynamical Behavior
Both climate science and control engineering are concerned with how systems respond to perturbations—but the questions they ask are shaped by the nature of their systems. In control engineering, stability analysis assesses whether a system returns to an equilibrium state after a disturbance. Concepts like asymptotic stability, bounded-input bounded-output stability, and Lyapunov functions are used to evaluate how reliably systems settle back into desired behavior [188, 189, 190, 191, 148, 192].
In contrast, climate science deals with systems that may not have a single stable equilibrium—or whose equilibria may themselves evolve over time due to external forcing. Rather than expecting the climate system to return automatically to a pre-disturbance state, the focus is on resilience: the ability of the system to absorb shocks without undergoing a qualitative, often irreversible shift. Tipping points, such as the collapse of the Atlantic Meridional Overturning Circulation or the rapid loss of the Greenland Ice Sheet, represent loss of resilience. After crossing such thresholds, returning to the original state may become physically impossible or require disproportionately large interventions [193, 194, 195, 196, 197, 198].
Several core differences emerge between the two fields:
  • Equilibria vs. Regimes: Control theory often studies stability around specific equilibria or trajectories. Climate science, instead, focuses on maintaining the system within broad functional regimes, without assuming an easy return to a specific baseline.
  • Perturbation Size: In engineering, stability typically concerns relatively small perturbations. In climate science, perturbations (e.g., anthropogenic CO2) can be large, pushing the system into fundamentally different basins of attraction.
  • Stability Tools: While both fields use phase space methods, climate science relies heavily on bifurcation diagrams, resilience metrics, early warning indicators (e.g., critical slowing down), and potential well analyses to understand the proximity to tipping points.
  • Control Objective: In control, stability ensures that desired performance is robust against disturbances. In climate science, interventions (e.g., emissions reductions, geoengineering) aim to maintain the climate system within acceptable safe boundaries, recognizing that recovery to the original state may not be possible.
Despite these differences, shared mathematical tools enable fruitful cross-fertilization. Concepts such as basins of attraction, stability margins, and feedback sensitivity provide a bridge to reason about climate resilience with a control-theoretic mindset. Moreover, understanding the climate system as a high-dimensional, feedback-rich dynamical system invites deeper adoption of dynamical systems theory methods, such as identifying early warning signals for critical transitions, optimizing interventions to maximize system resilience, and framing climate stability problems as control challenges under extreme uncertainty.
VII-F Model Calibration & Learning vs. System Identification
The problem of learning dynamic models from data and calibrating them to better reflect reality arises in both climate science and control engineering, though with different traditions and terminologies. However, the underlying ideas, identifying model parameters, uncovering dynamics, and improving predictive skill, are deeply paralleled across the two fields.
In control engineering, system identification refers to the process of building mathematical models of dynamic systems from observed input-output data [199, 200, 201, 202, 203, 204]. Techniques range from linear regression models for simple dynamics to nonlinear black-box models like neural networks, depending on the system complexity and the quality of available data. The primary objective is to either identify the full system dynamics (structure and parameters) or estimate unknown parameters within a known structure. Classical techniques include prediction error methods (PEM), subspace identification, and more recently ML-based methods for nonlinear systems. Rich toolboxes exist for these tasks, emphasizing guarantees like consistency, efficiency, and convergence.
Similarly, in climate science model calibration and learning are critical tasks. Climate models—whether EBMs, intermediate-complexity models, or full Earth system models (such as in Tab. III)—rely on parameterizations of unresolved or poorly understood processes (e.g., cloud dynamics, ocean eddies). Calibration involves tuning these parameters based on observational data, reanalysis products, or high-resolution simulations. Traditionally, this was done manually by expert judgment, but more systematic approaches like ensemble Kalman inversion, Bayesian calibration, and ML surrogates have become prominent [205, 206, 207, 208]. A key parallel between the two fields is that both deal with models that are incomplete, noisy, and computationally expensive. Learning hence must balance fidelity to data with physical realism. Several important connections and trends are worth highlighting:
  • Parameter Estimation: In control, unknown parameters are estimated through system identification. In climate modeling, parameters governing cloud microphysics, land-surface interactions, and biogeochemical cycles are similarly estimated using data assimilation, optimization, or statistical learning techniques.
  • Learning Full Models: Control engineering has long explored data-driven model discovery, from AutoRegressive with eXogenous inputs (ARX) models to modern neural network-based nonlinear models. In climate science, recent efforts have extended beyond tuning parameters to learning full sub-grid scale parameterizations (e.g., for convection, ocean eddies) using deep learning [209, 210]. Offline and online learning paradigms, as detailed in [210], now play a major role.
  • Hybrid Modeling: A growing trend in both fields is hybrid modeling, where physics-based models are augmented by data-driven components. In control, this parallels gray-box identification [211, 212, 213]; in climate studies, it includes neural-network parameterizations embedded within Earth system models (e.g., NeuralGCM [214]).
  • Challenges of Generalization and Stability: In both fields, learned models often face challenges when applied beyond the training regime (out-of-sample climates, unseen operating conditions). Issues such as physical constraint violations, instabilities, and error accumulation have motivated the development of physics-informed learning, regularized identification, and hybrid methods.
Recent landmark contributions in the climate community highlight the increasing adoption of machine learning:
  • ML techniques have been used for parameter tuning (e.g., Gaussian processes, ensemble Kalman inversions) [207].
  • Offline learning of parameterizations from high-resolution simulations to substitute for handcrafted physical models [215, 216].
  • Online learning of parameterizations during coupled model simulation to improve system statistics [217, 209].
  • Equation discovery using symbolic regression and sparse identification methods to find interpretable models for processes like ocean mesoscale eddies [218, 216].
Similarly, in dynamic systems, advances in learning dynamics and system identification under uncertainty, high-dimensionality, and partial observability mirror these developments. Techniques like Koopman operator theory, dynamic mode decomposition, and hybrid neural-physics models are increasingly central to the literature.
In short, the rich traditions of model learning and system identification in control engineering offer valuable insights for modern climate modeling. Conversely, the massive scale and complexity of climate applications are pushing the boundaries of what learning-based modeling needs to achieve—generalizability, physical consistency, and uncertainty-aware predictions—which can in turn inspire new developments in control and system identification methodologies.
VIII Tutorial Summary and Moving Forward
In this tutorial, we have produced the very first control systems-focused tutorial on climate science models and contemporary research problems in this rapidly evolving field. We hope that this tutorial bridges the disciplines of control engineering and climate science by exploring how control-theoretic principles can be applied to the behavior of climate systems. By viewing climate dynamics through a control systems lens, we highlight potential benefits for modeling, prediction, and strategic intervention. This perspective offers an alternative way to think about climate behavior and supports the idea of developing robust, physics-informed approaches for climate mitigation and adaptation. Integrating insights from both fields holds promise for better decision-making and more effective policy in response to climate change.
Building on this tutorial, we include ideas for future research papers that are well-contained in their scope.
  • Reachability Analysis: Given the following: (i) an ensemble of extremely complex, nonlinear, and large-scale climate models, (ii) some measurable initial conditions of climate states (or at least a region of initial climate state space), (iii) a long-term time-horizon of decades, and (iv) specific climate change mitigation strategies and various policy scenarios (as ones outlined in Fig. 9), one very relevant control engineering study could focus on investigating reachability analysis that considers (i)(iv). This results in realizing whether certain state space regions of climate spaces are indeed possible, under the various sources of uncertainty, without requiring running the large-scale models for different realizations of uncertainty. This problem of reachability is indeed well-contained if studied for one specific climate model as (8) for example. This problem is still theoretically challenging as reachability analysis for nonlinear systems is still an active field of research in systems science. Applying common reachability analysis in this context is of interest, but also designing specific climate science-based reachability analysis tools could be more pertinent.
  • Nonlinear Model Order Reduction: This augments or replaces the need to run expensive simulations that (ironically) consume large amount of electric power and hence further contribute to GHG emissions. As has been demonstrated in some recent work, the carbon footprint [219] of running climate models has exponentially increased recently. There has also been some proposed work to reduce the computational burden and the corresponding carbon footprint [220]. As a control theoretic alternative, model order reduction (MOR) for nonlinear dynamic systems can potentially be proposed to preserve the system properties (controllability, observability, stability) for the reduce order models. MOR algorithms basically reduce the state-space dimension, thereby resulting in faster forward simulations of the state-space dynamics. There has been recent significant advances in MOR theory for nonlinear systems [221, 222], and their investigation and application for large-scale complex climate climate models is a promising area of research.
  • Comparative analysis of system identification algorithms for climate models: There are virtually hundreds of control theoretic studies on the topic of system identification and learning dynamic models. We have avoided thoroughly discussing or citing any of this literature for brevity. Relevant to this paper, we note that it is still virtually unclear what is the overwhelmingly better method to perform large-scale system identification in complex nonlinear dynamic networks. In short, there is also an absence of computational studies that focus on comparing various identification and model learning approaches. Unfortunately, the abundant control engineering literature focuses on testing theoretical algorithms on toy examples that are not always meaningful or scalable. Dynamic climate models could hence be good benchmarks for testing the performance of system identification algorithms seeing that these models are riddled with uncertainty, extremely large-scale, often under-sensed, and have various forms of complex nonlinear relationships present in their dynamics.
  • Applications to Infrastructure Systems: Climate models enable predictive control opportunities in infrastructure systems affected by environmental variability. For instance in the water sector, projected shifts in precipitation, snowmelt, and drought severity can inform dynamic scheduling of water delivery, pump control, and reservoir operations to ensure supply reliability and minimize energy use—these have all been formulated and solved as optimal control problems in the literature. Similarly, in power systems, long-range and short-term climate forecasts can be integrated into real-time dispatch and control of renewable resources where temperature, wind, and solar radiation forecasts affect both load demand and generation profiles. These applications highlight the potential for closed-loop climate-informed control frameworks to enhance the resilience of critical infrastructure.
In the past two years, this tutorial presented itself as an opportunity for us to study and learn about a new problem, outside of the scope of our research interests. We were careful in our discussions and presented expositions. We used descriptive and welcoming language that can be understood by all controls and climate research readerships. We tried to not overhype the promise of control engineering in potentially addressing some climate science related problems, but also showcased promise of controls-guided climate science. We welcome suggestions, corrections, and clarifications from the curious and interested readers.